### v1

# AdvRepair: Provable Repair of Adversarial Attack

Abstract
Deep neural networks (DNNs) are increasingly deployed in safety-critical domains, but their vulnerability to adversarial attacks poses serious safety risks. Existing neuron-level methods using limited data lack efficacy in fixing adversaries due to the inherent complexity of adversarial attack mechanisms, while adversarial training, leveraging a large number of adversarial samples to enhance robustness, lacks provability. In this paper, we propose AdvRepair, a novel approach for provable repair of adversarial attacks using limited data. By utilizing formal verification, AdvRepair constructs patch modules that, when integrated with the original network, deliver provable and specialized repairs within the robustness neighborhood. Additionally, our approach incorporates a heuristic mechanism for assigning patch modules, allowing this defense against adversarial attacks to generalize to other inputs. AdvRepair demonstrates superior efficiency, scalability and repair success rate. Different from existing DNN repair methods, our repair can generalize to general inputs, thereby improving the robustness of the neural network globally, which indicates a significant breakthrough in the generalization capability of AdvRepair.
1 Introduction
In the last few years, deep neural networks (DNNs) have made significant advancements in various domains, including computer vision [3], natural language processing [16], and speech recognition [71]. However, the adoption of DNNs in safety-critical domains has been slow due to concerns regarding their dependability. One major concern is the vulnerability of DNNs to adversarial attacks [43, 14], where adversaries can manipulate input data in a way that is imperceptible to humans but can cause the model to make incorrect decisions. This vulnerability poses serious safety risks in applications such as autonomous vehicles [8] and medical diagnosis [68]. Therefore, it is crucial to develop and deploy DNN models with safety and reliability requirements in mind to fully realize their potential in safety-critical domains.
The repair of DNNs involves rectifying errors in the network by modifying its architecture or parameters. These errors can arise from various sources, including training errors [10], adversarial attacks [26, 13], backdoor attacks [40], and distribution drift [78, 62]. Regular and efficient repair during DNN deployment is crucial to mitigate the risks associated with these errors and ensure the performance and reliability of the network in practical applications. Traditional machine learning approaches for DNN repair include techniques such as adversarial training [24, 23, 65], input sanitization, fine-tuning, transfer learning [15, 81], and data augmentation [82, 51, 42]. These methods rely on a large amount of data and may exhibit suboptimal performance in situations with limited or low-quality data. Neuron-level fault localization and repair methods [61] aim to improve the accuracy and reliability of DNN repair by identifying and correcting errors at the individual neuron level. By leveraging these neuron-level techniques, researchers and practitioners can achieve more precise and reliable DNN repair with limited data.
At present, the efficacy of neuron-level repair methods utilizing only a limited number of samples for robust defense against adversarial attacks remains limited. These methods demonstrate effectiveness primarily in the context of backdoor attacks or individual safety properties. This limitation arises from the inherent complexity of adversarial attack mechanisms, which often exhibit a higher degree of intricacy compared to backdoor attacks. Moreover, diverse adversarial attacks targeting distinct samples correspond to a unique dense mixture of features [1], rendering it arduous to generalize parameter adjustments at the neuron level based on a small sample set. Furthermore, while techniques such as adversarial training leverage a large number of adversarial samples to bolster the robustness of neural networks, their efficacy lacks provability. Consequently, neural networks repaired using such techniques continue to harbor latent vulnerabilities in terms of robustness.
Our proposed methodology aims to address the issue of adversarial attacks by leveraging a limited set of adversarial samples. Specifically, our approach focuses on rectifying these erroneous samples while simultaneously maximizing accuracy and ensuring local robustness on these samples. Moreover, we aim to achieve this local robustness in a manner that generalizes across the entire dataset. To achieve these objectives, our method integrates an external indicator, a hard-coded characteristic function, which facilitates the allocation of patch modules in conjunction with the original network. This indicator identifies inputs within the local neighborhood of a specific sample and assigns the appropriate patch modules for repair. Additionally, for inputs outside the local neighborhood, a heuristic allocation mechanism is employed to enable generalization across the entire dataset.
Each patch module constitutes a fully connected neural network structure. The core concept of our approach lies in the utilization of reachability analysis through a linear approximation for the provable training of patch modules. Specifically, we employ the verification method DeepPoly [58] to establish a linear approximation of the output neurons. This approximation is leveraged to determine the distance between the target behavior and the current behavior, serving as the loss function. The patch modules are then trained to minimize this distance. Once the loss function reaches zero, the patch modules provide provable repairs for adversarial attacks within the perturbation region. By employing an external indicator to identify faulty inputs, adversarial attacks within the same perturbation region can be repaired using the same patch module. Inputs outside the perturbation range of limited adversarial samples are assigned by the indicator, leveraging a heuristic allocation mechanism, thereby ensuring the generalization of the repair. Through the comprehensive framework of our method, we effectively enhance the resilience of the network against adversarial attacks while preserving accuracy. By leveraging reachability analysis and patch modules, we provide provable repairs for adversarial samples, thereby improving the robustness of the network.
The contributions of this paper can be summarized as follows:
  • We propose a novel approach, named AdvRepair, specifically designed for provable repair of adversarial attacks using limited data. This work represents a significant breakthrough in achieving provable repair for adversarial robustness. Our method introduces patch modules that are seamlessly integrated into the original network architecture. By combining a loss function derived from formal methods, we train these patch modules to provide provable repairs for local robustness. Through a heuristic allocation mechanism, our approach successfully achieves the generalization of local robustness repair.
  • To address the efficiency challenges associated with formal verification in large-scale DNNs, we leverage patch modules to perform repairs at the feature space of the networks. This strategy enables our method to scale effectively across various network architectures, ensuring efficient and practical implementation.
  • We conduct a comprehensive evaluation of AdvRepair on three diverse datasets and multiple DNN architectures. Through extensive comparisons with state-of-the-art repair and adversarial training approaches, our method consistently demonstrates superior efficiency and scalability. Moreover, it exhibits remarkable generalization capabilities to handle general inputs, thereby significantly enhancing the overall robustness of the network. This advancement in generalization capabilities represents a substantial contribution of our method.
Overall, our proposed approach, AdvRepair, offers a pioneering solution for provable repair of adversarial attacks using limited data. It not only addresses the challenges posed by formal verification in large-scale DNNs but also achieves remarkable generalization capabilities, leading to improved network robustness.
2 Preliminary
In this section we recall some basic notations of DNN repair.
A deep neural network is a function ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$ that maps an input $x\in\mathbb{R}^{n_{0}}$ to an output $y\in\mathbb{R}^{n_{L}}$. We usually visualize a DNN $N$ as a sequence of $L$ layers, where the $i$th layer contains $n_{i}$ neurons each representing a real variable. Between two adjacent layers is typically a composition of an affine function and a non-linear activation function, and the DNN $N$ is the composition of the functions between layers. In many applications, DNNs are serving for classification tasks. In such a classification DNN ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$, every output dimension corresponds to a classification label, and the one with the maximum output value is the classification results that the DNN $N$ gives, i.e., $C_{N}(x)=\arg\max_{1\leq i\leq n_{L}}N(x)_{i}$, where $N(x)_{i}$ is the $i$th entry of the vector $N(x)$.
The notion of safety properties pertains to assertions that guarantee the absence of undesirable behavior. Within the context of DNNs, a safety property demands that a DNN operates correctly within a specified input range. The local robustness of a DNN refers to its ability to maintain stable and consistent predictions in the vicinity of its training data points, even in the presence of small perturbations or variations in the input. Here a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius. Formally, a local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, i.e., $\forall x\in B(x_{0},r),C_{N}(x)=C_{N}(x_{0})$. We denote this local robustness property as $(N,B(x_{0},r))$; it is obviously a safety property. An input $x^{*}\in B(x_{0},r)$ is an adversarial sample of the DNN $N$, if $C_{N}(x^{*})\neq C_{N}(x_{0})$, i.e., $x^{*}$ is a counterexample which violates the local robustness property $(N,B(x_{0},r))$.
In this work, we focus on the problem of repairing adversarial attacks with limited adversarial samples. Different from repairing backdoor attacks, not only does the buggy behavior of the given adversarial samples need fixing, but we also require that there should be no adversarial attacks around given adversarial samples, and that this enhancement of local robustness should generalize to samples across the whole dataset. Now we formally state the problem of repairing adversarial attacks as follows:
Given a DNN $N$ and a set of adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$, where $n$ can be significantly smaller than the size of the dataset, and each $x_{i}^{*}$ is obtained by an adversarial attack on an input $x_{i}$ with a given radius $r$, we need to construct a DNN $F$ which is locally robust on every $B(x_{i},r)$ while the accuracy is maintained and local robustness of other inputs is potentially improved.
3 Methodology
We focus on fixing adversarial attacks with limited data in this work, and the aims of the repair are at least threefold: For a given radius $r>0$,
  • the buggy behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ are fixed,
  • the DNN is locally robust in $B(x_{i},r)$ for $1\leq i\leq n$,
  • the accuracy of the DNN is maintained, and for as many input $x$ in the dataset as possible, the DNN is locally robust in $B(x,r)$.
Here the first objective is the fundamental requirement showing that the repair is successful on the given adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$. The second objective aims to ensure the local generalizability of the repair, whereby the absence of adversarial examples is guaranteed within the vicinity of the generated adversarial examples, while the third objective focuses on the global generalizability of the repair, wherein it simultaneously improves the robustness against all potential inputs in the dataset without significantly compromising the accuracy.
The current neural network repair methods under limited data conditions primarily involve parameter adjustments at the neuron level. While these methods demonstrate high efficacy against straightforward backdoor attacks, they face significant challenges in effectively addressing complex adversarial attacks. Even if the erroneous behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ are successfully repaired, such repairs lack the ability to generalize either locally or globally. Inspired by Reassure [22], we propose a patch-based repair method. A patch refers to a specific modification or alteration made to a software system or codebase; it is a discrete set of changes applied to fix a bug, enhance functionality, or address security vulnerabilities. Patch-based DNN repair involves the integration of an external indicator designed to identify buggy inputs, followed by the application of specialized patch modules to repair input sets which have similar behaviors. Each indicator here corresponds precisely to a robustness neighborhood $B(x_{i},r)$ that requires repair, and we can leverage verification tools to ensure that the repair within each robustness neighborhood is provable. Furthermore, for other inputs in the dataset, we can heuristically match patches that maximize their robustness, thereby endowing this repair approach with global generalization. In this section, we propose a patch-based repair method named AdvRepair to defend from adversarial attacks with limited data.
3.1 Structure of the repaired DNN
The main produce of the repaired network is depicted in Fig. 1. For an input $x\in\mathbb{R}^{n_{0}}$, the role of the indicator is to select an appropriate patch that, when applied, yields the sum of the output of the patch module and the output of the original DNN for $x$. This sum represents the output obtained after the repair process.
Figure 1: Framework of AdvRepair
The definition of an indicator function is standard:
Definition 1
Let $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties. The indicator function $\mathcal{I}_{\mathcal{C}}\colon\mathbb{R}^{n_{0}}\to\{0,1\}^{m}$ outputs in the $j$th entry as
$\mathcal{I}_{\mathcal{C}}(x)_{j}=\left\{\begin{array}[]{ll}1,&\text{ if }x\models X_{j},\\ 0,&\text{ otherwise},\end{array}\right.$
where $j=1,\ldots,m$.
Upon classification by the indicator, a set of patch modules is deployed to perform the repair. Here each patch module is specifically tailored to address an input set with the same local robustness property. The implementation of a patch module is a fully connected neural network in this work. For an input $x$ that is not within the input set of any indicator, i.e., $x\notin\bigcup_{i=1}^{n}B(x_{i},r)$, it cannot be identified by the indicators, but there is still potential adversarial attack in $B(x,r)$. In this situation, we heuristically allocate a patch module to this input to defend from adversarial attacks. Formally, the structure of the repaired DNN is as follows:
Definition 2
A repaired DNN is a tuple $F=(N,\mathcal{C},\mathcal{P},\tau)$, where
  • $N\colon\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L}}$ is the original DNN,
  • $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties,
  • $\mathcal{P}=(P_{1},\ldots,P_{m})^{\mathrm{T}}$ is a finite sequence of patch modules, each of which is a fully connected neural network,
  • and $\tau\colon\mathbb{R}^{n_{0}}\setminus\bigcup_{i=1}^{n}B(x_{i},r)\to 2^{\{1,\ldots,m\}}$ is a patch allocation function.
The semantics of the repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$ is a function
$F\colon\mathbb{R}^{n_{0}}$ $\to\mathbb{R}^{n_{L}},$
$x$$\mapsto\left\{\begin{array}[]{ll}\mathcal{I}_{\mathcal{C}}(x)^{\mathrm{T}}\mathcal{P}(x)+N(x),&\text{ if }x\in\bigcup_{i=1}^{n}B(x_{i},r),\\ \sum_{j\in\tau(x)}P_{j}(x)+N(x),&\text{ otherwise},\end{array}\right.$
where $\mathcal{P}(x)=(P_{1}(x),\ldots,P_{m}(x))^{\mathrm{T}}$.
Generally there are three ways to determine the number of the patch modules and define the indicator:
  • Property-guided. The property-guided way sets $\mathcal{C}_{\mathrm{P}}=\{B(x_{i},r)\mid i=1,\ldots,n\}$, i.e., the indicator judges which robustness region the input belongs to, and the corresponding patch module defends against adversarial attacks in this specific region.
  • Untargeted label. A simple label-guided way takes every classification label as an input property. i.e., $\mathcal{C}_{\mathrm{L}}=\{1,\ldots,n_{L}\}$, where $x\models\ell$ iff $C_{N}(x)=\ell$. Assuming that $x$ is correctly classified by $N$, the patch module to defend against adversarial attacks in $B(x,r)$ is determined by $C_{N}(x)$, and it works for fixing potential adversarial attacks of all the robustness regions where the ground truth is $C_{N}(x)$.
  • Targeted label. We consider the set of label pairs $\mathcal{C}_{\mathrm{L}}=\{(\ell_{0},\ell_{1})\in\{1,\ldots,n_{L}\}^{2}\mid\ell_{0}\neq\ell_{1}\}$, where $x\models(\ell_{0},\ell_{1})$ iff $N(x)_{\ell_{0}}>N(x)_{\ell_{1}}>N(x)_{\ell}$ for all $\ell\neq\ell_{0},\ell_{1}$, i.e., the indicator determines which label $\ell_{1}$ is the most likely when attacking the ground truth $\ell_{0}$, assuming that $x$ is correctly classified by $N$, and the corresponding patch module fixes adversarial attacks of this specific pattern.
The property-guided way has advantages when the number $n$ of the given adversarial samples is small, while the label-guided ways work well in the situation when the number $n_{L}$ of the output labels is small. For the label-guided ways, the patch allocation function $\tau$ can be simply defined as the corresponding indicator, since this indicator can naturally generalize to other input. In the following, we present our approach in the property-guided way, and it can be easily adapted to the label-guided ways.
3.2 Training the patch modules
The first challenge is to train the patch modules, which are each implemented as a fully connected neural network, to repair each robustness property $(N,B(x_{i},r))$ with a provable guarantee. To achieve provability, we employ formal verification in training these patch modules. The result of verification on a robustness property can be transformed into a loss function; once this loss reaches zero, the property is verified to be true, and the patch module in this stage can defend against all possible adversarial attacks in this robustness region. This idea has been widely adopted in DNN repair [21, 39].
In this work, we employ DeepPoly [58] as the verification engine. DeepPoly uses abstract interpretation to give every neuron an upper/lower bound in the form of an affine function, where only variables in previous layers occur, and the numerical bound can be derived by propagating backwards these affine upper/lower bounds to the input layer. Recall that a local robustness property $(N,B(x_{i},r))$ holds iff for any $x\in B(x_{i},r)$ and $\ell\neq\ell_{0}=C_{N}(x_{i})$, $N(x)_{\ell}<N(x)_{\ell_{0}}$, so DeepPoly calculates the abstraction of $N$ on $B(x_{i},r)$ for every neuron and the expressions $N(x)_{\ell}-N(x)_{\ell_{0}}$ for $\ell\neq\ell_{0}$. From the abstraction, we can obtain an affine function in which there are only input variables as a sound upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$, i.e.,
$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},$ (1)
where $\alpha_{\ell}\in\mathbb{R}^{n_{0}}$ and $\beta_{\ell}\in\mathbb{R}$ are constants. It is easy to obtain the numerical upper bound of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on the box region $B(x_{i},r)$. If this upper bound is negative for every $\ell\neq\ell_{0}$, then the local robustness property $(N,B(x_{i},r))$ is verified to be true by DeepPoly. Therefore, it is natural to use this upper bound as the loss function to train the corresponding patch module.
Definition 3
For a local robustness property $\varphi=(N,B(x_{i},r))$, we define the safety violated loss function of $\varphi$ as
$\mathcal{L}(\varphi)(x)=\sum_{\ell\neq\ell_{0}}\max(\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},0),$
where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is the upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$ given by DeepPoly, as shown in Eq. (1). For local robustness properties $\varphi_{1},\ldots,\varphi_{k}$ which share the same ground truth label $\ell_{0}$, we define $\mathcal{L}(\bigwedge_{i=1}^{k}\varphi_{i})=\sum_{i=1}^{k}\mathcal{L}(\varphi_{i})$.
For a local robustness property $\varphi=(N,B(x_{i},r))$, its safety violated loss function $\mathcal{L}(\varphi)$ being 00 implying that $\varphi$ is verified to be true by DeepPoly.
Theorem 3.1
Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
When we train a patch module $P$ on the DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$, the maximum of the safety violated loss function of a local robustness property $\varphi$, namely $\mathcal{L}^{*}(\varphi)$, is parameterized with the weights $w$ in $P$, and this $\mathcal{L}^{*}(\varphi)(w)$ is the loss function for training this patch $P$. By employing the standard gradient descent method, we can train the patch $P$, until the loss reaches 00 or the number of epochs reaches a threshold $R$. We claim that, even if $\mathcal{L}^{*}(\varphi)(w)$ does not reach 00, the repair does not necessarily fail, since the upper bound $\mathcal{L}^{*}(\varphi)$ given by DeepPoly is not exact; as it decreases in the training process, $F$ is growing more likely to satisfy $\varphi$. Also, to improve the precision of $\mathcal{L}^{*}(\varphi)$ obtained from DeepPoly, we employ the input interval partitioning technique from ART [39]. It selects the partition dimension by computing the multiplication of the partial derivative of the safety violated loss function $\mathcal{L}(\varphi)(x)$ and the size of the input interval in the corresponding dimension, and bisects the box region over the dimension with the maximum score.
Algorithm 1 AdvRepair (Property-guided)
1:
2:Original DNN $N$, pairs of input and its adversarial examples $\{(x_{i},x_{i}^{*})\}_{i=1}^{n}$, and radius $r>0$
3:
4:A repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$
5:$\mathcal{C}\leftarrow(B(x_{1},r),\ldots,B(x_{n},r))$, $\mathrm{iter}\leftarrow 0$
6:$\mathrm{Init}(\mathcal{P})$triangleright Initialize the patch modules $\mathcal{P}=(P_{1},\ldots,P_{n})$
7:$D[\cdot,\cdot]\leftarrow\{\}$triangleright A dictionary where $D[i]$ stores properties fixed with $P_{i}$
8:for $i\leftarrow 1~{}\text{to}~{}n$ do
9:     $D[i]\leftarrow\{(F,B(x_{i},r))\}$
10:$E\leftarrow\{1,\ldots,m\}$triangleright Record the properties not fixed yet
11:while $\mathrm{iter}<M$ do $\triangleright$ $M$: Maximum number of iterations
12:     $\mathrm{iter}\leftarrow\mathrm{iter}+1$
13:     for $j\in E$ do
14:         $(P_{j},T,\mathrm{Repaired})\leftarrow\textsc{Train}(N,P_{j},D[j])$triangleright Alg. 2
15:         if $\mathrm{Repaired}=\mathbf{True}$ then
16:              $E\leftarrow E\setminus\{j\}$
17:         else
18:              for $\varphi\in T$ do $\triangleright$ We write $\varphi=(F,X)$
19:                  for $d\leftarrow 1~{}\text{to}~{}n_{0}$ do
20:                       $\operatorname{score}_{d}\leftarrow\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|$
21:                  $d^{*}\leftarrow\arg\max_{d}\operatorname{score}_{d}$
22:                  $X_{1},X_{2}\leftarrow\mathrm{Bisect}(X,d^{*})$triangleright Bisection on the $d^{*}$th dimension
23:                  $\varphi_{1}\leftarrow(F,X_{1})$, $\varphi_{2}\leftarrow(F,X_{2})$, $D[j]\leftarrow(D[j]\cup\{\varphi_{1},\varphi_{2}\})\setminus\{\varphi\}$
24:     if $E=\{\}$ then
25:         return $F$triangleright Repair with provable guarantee
26:return $F$triangleright Repair without provable guarantee
Algorithm 2 Patch training
1:
2:Original DNN $N$, DNN $P$ as a patch module, and a finite set $U$ of local robustness properties
3:
4:The optimized patch module $P$, the set $T\subseteq U$ of properties to be refined, and whether the properties in $U$ have all been fixed with a provable guarantee
5:function Train($N,P,U$)
6:     $\mathrm{epoch}\leftarrow 0$
7:     while $\mathrm{epoch}<R$  do $\triangleright$ $R$: maximum number of epochs
8:         $\mathrm{epoch}\leftarrow\mathrm{epoch}+1$
9:         $w\leftarrow w-\eta\cdot\nabla\mathcal{L}^{*}(\bigwedge U)(w)$triangleright $w$: the weights in $P$
10:         if $\mathcal{L}^{*}(\bigwedge U)(w)=0$ then $\triangleright$ $\mathcal{L}^{*}(\bigwedge U):=\mathcal{L}^{*}(\bigwedge_{\varphi\in U}\varphi)=\sum_{\varphi\in U}\mathcal{L}^{*}(\varphi)$
11:              return $(P,\{\},\mathbf{True})$
12:     $T\leftarrow\mathrm{Slice}_{K}(\mathrm{ArgSort}_{\varphi}\{\mathcal{L}^{*}(\varphi)(w)\mid\mathcal{L}^{*}(\varphi)(w)>0,\varphi\in U\})$
13:     return $(P,T,\mathbf{False})$
The main algorithm of AdvRepair is shown in Alg. 1. We construct a dictionary $D$, where $D[i]$, initialized as $\{(F,B(x_{i},r))\}$, stores the properties that the $i$th patch module $P_{i}$ repairs (Line 3–5, Alg. 1). The repair process has at most $M$ iterations. In each iteration, we train the patches which does not yet provide provable repair, i.e., those whose index is in $E$. The algorithm of training a specific patch module $P$ for repairing a set $U$ of properties is shown in Alg. 2, which outputs the optimized patch module $P$, the set $T$ of properties to be refined, and a boolean label recording whether the repair of the patch $P$ has a provable guarantee. A standard gradient descent procedure is run until a provable repair is achieved, i.e., $\mathcal{L}^{*}(\bigwedge U)(w)=0$, or the number of epochs reaches a threshold $R$ (Line 3–7, Alg. 2). After that, if it is still not provable, we sort $\mathcal{L}^{*}(\varphi)(w)$ for $\varphi\in U$ from the largest to the smallest, and extract the largest $K$ ones (whose value is strictly larger than 00) as the properties to be refined (Line 8, Alg. 2). For a property $\varphi=(F,X)$ to be refined, we select an input dimension to bisect the input space $X$. For an input dimension $d$, we define
$\mathrm{score}_{d}=\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|,$
where $\partial_{d}\mathcal{L}(\varphi)(x)$ is the partial derivative of $\mathcal{L}(\varphi)(x)$ on the $d$th dimension. We choose the dimension with the largest score to bisect $X$, and the property $\varphi=(F,X)$ is refined to two properties $\varphi_{1}=(F,X_{1})$ and $\varphi_{2}=(F,X_{2})$, recorded in the dictionary $D$ (Line 14–19, Alg. 1). At the end of each iteration, we check whether all the patches provide provable repair; if so, it terminates immediately and outputs the current repaired DNN $F$, and this repair is provable (Line 20–21, Alg. 1). If $E$ is still non-empty after $M$ iterations, it outputs $F$ without provable guarantee (Line 22, Alg. 1). Although we are focusing on fixing adversarial attacks in this work, AdvRepair also works for repair safety properties of DNNs by making minor changes in Alg. 1.
In AdvRepair, we follow a classical way of defining the loss function with the linear relaxation obtained from DeepPoly, and incorporate it into our patch-based repair framework. This combination has several advantages. First, every patch is responsible for repairing properties in a specific pattern, which makes it much easier to obtain a provable repair; even if the repair is not provable, the safety violated loss is significantly declining in the training process, and it is highly possible that the repaired DNN is locally robust. Different from the neuron-level repair, there is no modification on the original DNN $N$ in AdvRepair, and due to the design of the indicator, the patch modules does not affect the behaviors of other input before $\tau$ allocates a patch to it. This mechanism avoids the drawdown of accuracy, which is quite severe in many neuron-level repair methods. Also, by allocating patches with $\tau$, we can achieve good generalization of our repair to other inputs, and it is the key to solving this essential challenge in fixing adversarial attacks. In the following, we will present how to construct the patch allocation function $\tau$ to improve generalization.
3.3 Patch allocation
For the property-guided method in AdvRepair, we need to define a patch allocation function $\tau$ because the construction of patches in this way does not naturally generalize to other inputs. This patch allocation function $\tau$ is the key to achieving good generalization of the patch modules trained locally to global inputs.
To match the best patch modules for an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we need to extract some specific patterns of the patch modules which generalize to other inputs and their adversarial robustness. Such patterns may include:
  • the ground truth of $x^{\prime}$, which can be represented with $\ell_{0}=C_{N}(x^{\prime})$,
  • the targeted label, i.e., the index of the second largest entry in $N(x^{\prime})$,
[…]
In the implementation of $\tau$, the patches with the same ground truth label as $x^{\prime}$ is alternative for $x^{\prime}$, and we allocate the ten with the maximum cosine similarity between $N(x^{\prime})$ and $N(x_{i})$ to $x^{\prime}$. Here we do not match with the targeted label because it is possible that more than one labels can potentially attack $N$ in the robustness region $B(x^{\prime},r)$, and matching by the cosine similarity can cover this possibility.
We seemingly have problem when $x^{\prime}$ is not correctly classified by $N$, because in this case we may allocate inappropriate patches. If $x^{\prime}$ is an adversarial example, we can employ sampling-based methods like [70] to detect it and recognize its correct classification with a high probability. Otherwise, the wrong classification of $x^{\prime}$ may result from backdoor attacks, biased training data, overfitting, etc, and such situations are beyond the scope of fixing adversarial attacks in this work.
Once a set of patch modules has been allocated to an input, it works in the whole robustness region of this input. For an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we allocate it with $\tau(x^{\prime})$, a set of patch modules. After that, all the inputs $x\in B(x^{\prime},r)$ share the same set of patch modules as $x^{\prime}$, i.e., $F(x)=\sum_{j\in\tau(x^{\prime})}P_{j}(x)+N(x)$, instead of allocating $x$ with $\tau(x)$. In this sense, our defense is constructed before adversarial attacks.
3.4 Repair in a feature space
To repair an extremely large DNN with high input dimensionality, the memory and computational overheads associated with the linear relaxations obtained from formal verification are notably significant and, at times, prohibitive. Also, faced with a huge DNN, the precision of abstraction-based verification methods like DeepPoly is severely harmed during a large number of layer-by-layer propagation. Consequently, the key to adopting AdvRepair to large DNNs is to reduce the input dimensionality and the size of the neural network in formal verification.
In the popular architectures of convolutional neural networks, the convolutional layers play the role of extracting important features from data, followed by several fully connected layers for classification according to these extracted features, so we call such a fully connected layer a feature space.
Definition 4
Let $N\colon\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L}}$ be a DNN where the function from the $i$th layer to the $(i+1)$th is $f_{i}$, i.e., $N=f_{L-1}\circ\cdots\circ f_{0}$. For $l\in[1,L]$, the feature space of the $l$th layer is $\mathbb{R}^{n_{l}}$, and the behavior of the input $x\in\mathbb{R}^{n_{0}}$ in this feature space is $N_{0,l}:=f_{l-1}\circ\cdots\circ f_{0}(x)$. The neural network starting with the feature space of the $l$th layer in $N$ is $N_{l,L}=f_{L-1}\circ\cdots\circ f_{l}$.
The dimensionality of a feature space is usually far smaller than that of the input layer. If we start with a feature space in employing DeepPoly, it does not need to calculate the abstraction of a great deal of neurons in the convolutional layers, thus getting its efficiency and precision enhanced.
To conduct repair in a given feature space with AdvRepair, we first give an approximation of the feature space on every $B(x_{i},r)$. Here we do not require that this approximation should be a sound abstraction of the real semantics of the feature space on $B(x_{i},r)$, because calculating such an over-approximation with high precision is quite time-consuming. Instead, we simply sample in $B(x_{i},r)$ and obtain the tightest box region that contains the behaviors of the samples in the feature space as the approximation. This approximation is not sound yet, but it does not mean that the inputs in $B(x_{i},r)$ whose behaviors in the feature space are not within this approximation will not be repaired by the corresponding patch module $P_{i}$, because the indicator $\mathcal{I}_{\mathcal{C}}$ still remains the same. Since the approximation of the feature space is not sound, the repair does not have a provable guarantee. It is also worth mentioning that, even if two properties with different ground truth labels have their feature spaces overlapping, our repair still works. In this situation, although two contradicted properties are involved in the training process, but they must correspond to two different robustness regions $B(x_{i},r)$, and thus different patch modules $P_{i}$. The correspondence of the robustness regions $B(x_{i},r)$ and the patch $P_{i}$ is always preserved, so $P_{i}$ is always working for repairing on $B(x_{i},r)$ with the correct classification label $\ell_{0}$.
After we obtain an approximation $X_{(i)}$ for each $B(x_{i},r)$ in the feature space, we employ Alg. 1, where the initialization of $D[i]$ in Line 5 is $(N_{l,L}+P_{i},X_{(i)})$ instead. Since the repair in the feature space is not provable, the safety violation loss function $\mathcal{L}(\varphi)$ for $\varphi=(N_{l,L}+P_{i},X_{(i)})$ in Def. 3 is modified to be
$\mathcal{L}(\varphi)(x)=\sum_{\ell\neq\ell_{0}}(\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}),$
where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ for $\ell\neq\ell_{0}$ is obtained by DeepPoly abstracting $N_{l,L}+P_{i}$ on $X_{(i)}$, and $\mathcal{L}^{*}(\varphi)$ is modified accordingly. Also, we do not run Line 6–7 in Alg. 2 and skip Line 6, 11–12 and 20–21 in Alg. 1, so that the loss $\mathcal{L}^{*}$ decreases as much as possible and we can achieve a better repair performance.
Since we are here considering the feature space in AdvRepair, it is natural to take it into consideration the cosine similarity and the distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space. If the ground truth label of $x^{\prime}$ and $x_{i}$ matches, a high cosine similarity and a small distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space strongly indicate that the patch module $P_{i}$ for $x_{i}$ is effective in fixing adversarial attacks in $B(x^{\prime},r)$. In this case, the patches with the same ground truth label as $x^{\prime}$ are still alternative for $x^{\prime}$, but we allocate the ten patches with the maximum value of
$\frac{\langle N_{0,l}(x^{\prime}),N_{0,l}(x_{i})\rangle}{\|N_{0,l}(x^{\prime})\|_{2}\cdot\|N_{0,l}(x_{i})\|_{2}}\cdot\|N_{0,l}(x^{\prime})-N_{0,l}(x_{i})\|_{2}.$
Repair in a feature space is an effective way to make AdvRepair scale on large DNNs. Although we sacrifice provability, its repair performance and generalization are extremely ideal in practical scenarios, which we will see in our experimental evaluation.
4 Experimental Evaluation
In this section, we evaluate AdvRepair by answering the following research questions:
RQ1:What is the overall performance of AdvRepair in repairing local robustness and correcting safety property violations?
RQ2: Does the repaired DNN exhibit the capability to defend against new adversarial attacks?
RQ3: Does AdvRepair have scalability to repair large networks?
RQ4: What impact does the number of adversarial samples needed for repair have on the effectiveness of the repair process?
4.1 Setup
All the experiments are conducted on a machine equipped with an AMD EPYC 7763 64-Core Processor, 256 GB of memory, and an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory. Each experiment has a timeout set to 10 000 seconds.  [^1] AdvRepair is available at an anonymous link: https://drive.google.com/drive/folders/1NU170UZ0XmYKgg6diKZc-XMcqsUAVEib?usp=drive˙link
Dataset. We conduct evaluations of our approach on three datasets: MNIST [35], CIFAR-10 [34] and ACAS Xu [69, 29]. On MNIST, we train two fully connected neural networks and one convolutional neural network. On CIFAR-10, we train a VGG19 [54] and a ResNet18 [20]. We assess our approach across 35 ACAS Xu DNNs, which comprises six hidden layers, each housing 50 neurons. As documented in [32], these DNNs are expected to satisfy Property-2 but exhibit violations. Detailed model structures and hyperparameters are in Appendix 0.C. For the local robustness repair task, we subject the DNNs trained on MNIST and CIFAR-10 to PGD [43], generating 50, 100, 200, 500 and 1 000 adversarial samples. For MNIST, the radius $r$ is set to be 0.05, 0.1 and 0.3, while for CIFAR-10 it is set to be $\frac{4}{255}$ and $\frac{8}{255}$. On ACAS Xu, our objective is to repair the violation of Property-2 while preserving the original performance of the ACAS Xu DNNs. We sample 500 counterexamples as the buggy inputs for repair.
Metrics. To assess generalization, we sample ten adversarial examples on each $B(x_{i},r)$ different from $x_{i}^{*}$ to form a generalization set $D_{\mathrm{g}}$ for MNIST and CIFAR-10, and use an independent set of 5 000 counterexamples for ACAS Xu. Besides these, we have an independent test set $D_{\mathrm{t}}$ of size 10 000, 10 000 and 5 000 for MNIST, CIFAR-10 and ACAS Xu, respectively. We employ AutoAttack [14] to attack the repaired DNN $F$ on $\{x_{i}\}_{i=1}^{n}$ and $D_{\mathrm{t}}$ with the same radius $r$. The metrics we use include repair success rate ($\mathrm{RSR}$), repair generalization rate ($\mathrm{RGR}$), drawdown ($\mathrm{DD}$), defense success rate ($\mathrm{DSR}$), and defense generalization success rate ($\mathrm{DGSR}$), defined as follows:
$\mathrm{RSR}$ =|$=\frac{|\{\,i\mid C_{F}(x_{i}^{*})=C_{N}(x_{i})\,\}|}{n},\qquad\mathrm{RGR}=\frac{|\{{\,x\in D_{\mathrm{g}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{g}}\right|},$
$\mathrm{DD}$ =|$=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{N}(x)=\ell_{x}\,\}|-|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{t}}\right|},$
$\mathrm{DSR}$ =|$=\frac{|\{\,i\mid\exists x\in\mathrm{AA}(F,B(x_{i},r)),C_{F}(x)\neq C_{N}(x_{i})\,\}|}{n},$
$\mathrm{DGSR}$ =|$=\frac{|\{\,x\in D_{\mathrm{t}}\mid\exists x^{\prime}\in\mathrm{AA}(F,B(x,r)),C_{F}(x^{\prime})\neq C_{N}(x)\,\}|}{\left|D_{\mathrm{t}}\right|},$
where $\ell_{x}$ denotes the ground truth of $x$, and $\mathrm{AA}(\varphi)$ represents the set of potential adversarial samples obtained by attacking the local robustness property $\varphi$ with AutoAttack. On ACAS Xu, the metric Drawdown is replaced with Fidelity Drawdown ($\mathrm{FDD}$) because there is no labelled ground truth for inputs on ACAS Xu:
$\mathrm{FDD}=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=C_{N}(x)\,\}|}{|D_{\mathrm{t}}|}.$
The metrics $\mathrm{RSR}$ and $\mathrm{DD}$ (or $\mathrm{FDD}$) evaluates the overall performance of DNN repair by measuring the percentage of buggy inputs successfully repaired and how much the accuracy decreases, while $\mathrm{RGR}$, $\mathrm{DSR}$ and $\mathrm{DGSR}$ reflects how the repair generalizes to the robustness regions $B(x_{i},r)$ and to other inputs.
Baselines. We compare AdvRepair with the state-of-the-art repair methods including CARE [61], APRNN [63], PRDNN [60], REASSURE [22] and ART [39]. Since APRNN and PRDNN require selecting a layer to repair on, we traverse all eligible layers across each baseline, and report the layer that exhibit the best performance. Additionally, for the local robustness repair task, we compare our approach with an adversarial training method TRADES [83], where we select the trained model with the best performance in 200 epochs with their default parameters.
Table 1: Results of repairing local robustness, where REASS and TRADE denote REASSURE and TRADES, respectively, and – means timeout or memory overflow.
Model $r$ $n$ RSR/DD/Time/s
CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
FNN_ small 0.05 50 8.0 100.0 100.0 100.0 54.0 10.0 100.0 1.0 30.9 19.7 0.0 3.3 86.8 0.0 2.0 35.9 3.5 381.3 5.2 257.7 14.3
100 2.0 100.0 100.0 100.0 2.0 10.0 100.0 0.1 25.8 29.6 0.0 86.8 87.0 0.0 2.3 80.8 9.6 762.5 13.4 318.5 24.8
200 4.5 100.0 0.0 100.0 3.5 11.0 100.0 0.4 39.5 66.2 0.0 86.8 86.5 0.0 3.2 268.2 82.8 1486.6 24.6 422.3 60.6
500 5.8 100.0 0.0 100.0 2.8 11.0 100.0 1.7 56.1 68.3 0.0 86.8 86.8 0.0 8.2 1358.4 73.0 3812.3 54.1 688.7 236.7
1000 11.4 100.0 0.0 3.9 11.0 100.0 16.6 43.1 69.2 86.8 87.0 0.0 17.3 3473.5 34.3 103.0 1247.5 756.7
0.1 50 2.0 100.0 100.0 100.0 8.0 14.0 100.0 0.0 71.3 50.9 0.0 86.8 86.5 0.0 2.2 51.6 3.8 358.8 8.2 393.0 15.5
100 2.0 100.0 0.0 100.0 13.0 9.0 100.0 0.0 62.4 48.5 0.0 86.8 87.0 0.0 1.7 92.2 72.8 737.7 13.4 321.4 37.3
200 1.0 100.0 0.0 100.0 10.5 12.0 100.0 0.3 60.1 63.9 0.0 86.8 86.5 0.0 2.8 321.8 149.2 1403.8 22.7 416.7 62.5
500 0.4 100.0 0.0 100.0 10.0 14.0 100.0 -0.1 51.6 71.8 0.0 86.8 86.8 0.0 5.1 1324.6 27.9 3482.9 53.7 564.4 193.2
1000 0.6 100.0 0.0 10.0 14.0 100.0 0.1 42.4 73.0 86.8 86.8 0.0 8.0 3844.8 33.9 104.5 1733.2 768.1
0.3 50 0.0 100.0 100.0 100.0 8.0 8.0 94.0 0.0 77.0 42.9 0.0 86.8 86.8 0.0 1.8 55.3 3.6 328.3 13.3 255.7 18.6
100 6.0 100.0 100.0 100.0 13.0 11.0 99.0 7.0 74.1 61.5 0.0 86.8 87.0 0.0 2.6 115.7 9.5 670.3 13.4 322.3 52.0
200 0.0 100.0 0.0 100.0 10.5 8.0 99.5 0.0 62.8 54.2 0.0 86.8 87.0 0.0 3.2 364.9 132.1 1306.2 23.2 405.6 115.8
500 0.0 100.0 0.0 100.0 10.0 10.0 99.4 -0.1 57.1 65.7 0.0 86.8 86.3 0.0 4.8 1934.4 147.9 3248.1 52.5 555.8 255.9
1000 0.0 100.0 0.0 9.8 11.0 99.6 0.0 58.3 83.9 86.8 86.8 0.0 7.4 3759.6 44.7 103.2 2116.1 935.8
FNN_ big 0.05 50 6.0 100.0 100.0 100.0 98.0 24.0 100.0 2.0 0.7 3.6 0.0 0.4 85.8 0.0 1.7 89.5 11.2 1807.6 5.7 436.5 14.5
100 10.0 100.0 100.0 100.0 78.0 9.0 100.0 2.8 0.8 16.0 0.0 -0.1 86.9 0.0 3.6 171.1 24.3 3673.7 14.3 503.5 29.1
200 10.0 100.0 100.0 100.0 73.0 6.0 100.0 4.8 1.0 28.1 0.0 0.1 88.3 0.0 4.5 419.0 92.3 7399.1 24.3 628.3 51.2
500 10.0 100.0 100.0 67.0 10.0 100.0 6.1 1.6 25.8 -0.1 87.1 0.0 9.2 1390.1 657.6 54.4 1080.2 290.6
1000 11.2 100.0 100.0 71.9 20.0 100.0 6.2 1.9 28.0 0.1 85.8 0.0 17.9 4139.0 5499.2 108.1 2167.6 1033.2
0.1 50 18.0 100.0 100.0 100.0 8.0 8.0 100.0 13.7 6.7 13.9 0.0 0.8 87.4 0.0 2.7 88.2 10.5 1612.1 7.1 442.2 13.9
100 2.0 100.0 100.0 100.0 27.0 11.0 100.0 0.1 5.6 25.8 0.0 8.8 87.4 0.0 2.9 173.7 22.8 3337.2 13.5 514.6 32.1
200 17.5 100.0 100.0 100.0 15.5 10.0 100.0 10.6 6.1 27.7 0.0 0.7 87.4 0.0 6.4 445.5 95.8 6801.2 23.3 639.2 75.4
500 0.6 100.0 100.0 15.4 15.0 100.0 1.3 7.4 33.7 0.5 88.3 0.0 6.3 1403.6 684.8 57.1 1048.6 356.8
1000 0.5 100.0 100.0 26.4 11.0 100.0 0.7 8.3 33.3 0.5 86.9 0.0 11.2 4321.6 4726.4 107.8 2148.3 735.9
0.3 50 0.0 100.0 100.0 100.0 0.0 14.0 100.0 0.0 27.5 33.3 0.0 0.5 85.8 0.0 1.8 119.5 11.6 1280.9 7.1 439.2 17.1
100 1.0 100.0 100.0 100.0 4.0 14.0 100.0 0.0 31.8 56.6 0.0 1.8 85.8 0.0 2.1 253.1 27.0 2612.3 13.5 518.9 37.0
200 1.0 100.0 100.0 100.0 2.0 10.0 100.0 0.2 28.5 28.4 0.0 0.4 86.9 0.0 3.7 630.5 105.6 5186.7 23.9 643.3 100.4
500 0.4 100.0 100.0 2.2 9.0 100.0 0.2 23.9 29.5 0.6 86.9 0.0 4.9 2117.7 671.8 55.2 1047.6 365.4
1000 0.6 100.0 100.0 4.8 12.0 100.0 0.7 23.6 26.3 0.8 85.8 0.0 10.9 5480.3 4297.3 108.8 2158.5 1246.4
CNN 0.05 50 0.0 100.0 100.0 100.0 100.0 0.0 100.0 2.4 0.8 0.0 0.0 0.9 88.5 0.0 3.5 6.2 57.9 509.3 4.3 789.9 14.1
100 0.0 100.0 100.0 100.0 92.0 60.0 100.0 2.4 0.7 0.1 0.0 0.2 88.5 0.0 6.0 5.1 99.0 910.7 10.3 879.0 23.4
200 0.0 100.0 100.0 100.0 88.0 0.0 100.0 2.7 1.0 0.0 0.0 0.0 88.5 0.0 9.8 19.0 182.9 1825.6 18.0 969.9 83.6
500 0.0 100.0 100.0 92.2 100.0 3.4 1.2 -0.1 0.0 0.0 26.9 146.2 547.8 40.4 221.9
1000 0.0 100.0 100.0 91.2 100.0 7.5 1.8 0.0 0.0 0.0 38.1 549.1 1611.6 80.9 1001.9
0.1 50 0.0 100.0 100.0 100.0 94.0 0.0 100.0 0.6 1.0 0.1 0.0 0.4 88.0 0.0 2.6 45.6 56.6 456.0 5.7 938.0 12.7
100 0.0 100.0 100.0 100.0 94.0 10.0 100.0 1.3 1.0 0.1 0.0 0.7 88.7 0.0 5.1 9.8 104.8 896.8 10.4 1022.8 29.5
200 0.0 100.0 100.0 100.0 91.5 20.0 100.0 2.8 2.2 0.3 0.0 0.4 88.5 0.0 8.1 41.1 178.5 1800.8 17.7 1108.1 75.1
500 0.0 100.0 100.0 91.6 100.0 9.1 1.7 0.4 0.5 0.0 24.9 139.2 458.5 40.6 337.0
1000 0.0 100.0 100.0 92.8 100.0 5.4 2.4 0.2 0.5 0.0 32.0 492.7 1526.0 80.4 700.2
0.3 50 0.0 100.0 100.0 100.0 0.0 10.0 100.0 0.0 43.2 7.2 0.0 0.8 88.5 0.0 2.2 3.8 59.9 453.8 5.6 808.2 15.0
100 0.0 100.0 100.0 100.0 0.0 10.0 100.0 0.0 38.8 5.0 0.0 0.2 89.3 0.0 2.6 29.4 105.4 907.5 10.1 899.5 32.6
200 0.0 100.0 100.0 100.0 0.0 0.0 100.0 0.0 36.8 3.6 0.0 0.2 89.3 0.0 4.6 59.9 196.2 1793.7 17.8 1014.7 87.9
500 0.0 100.0 100.0 0.2 100.0 0.5 48.8 3.2 0.3 0.0 9.2 167.9 518.8 39.9 361.4
1000 0.0 100.0 100.0 0.6 100.0 0.1 56.5 2.2 0.4 0.0 16.8 423.8 1667.0 80.4 1282.6
4.2 Repair performance
For the overall repair performance, we focus on $\mathrm{RSR}$, $\mathrm{DD}$ and running time, which indicate whether the buggy behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ can be fixed, whether the accuracy is maintained, and the efficiency, respectively.
The results of the local robustness repair task on MNIST are summarized in Table 1. In the experiments of repairing local robustness, PRDNN, REASSURE and AdvRepair all achieve a perfect 100% $\mathrm{RSR}$, attributed to their provable constructions. APRNN, while provable, exhibits limitations in effectively repairing specific instances within the FNN_small model. Despite being a provable method, ART is less effective in repairing tasks involving high-dimensional inputs. CARE and TRADES, which do not have provable guarantee, demonstrate lower $\mathrm{RSR}$ compared to the provable approaches. CARE exhibits suboptimal repair performance for all the three models, possibly due to the intricate relationship between local robustness repair and the numerous parameters in the DNNs. The challenge lies in resolving such errors through fault localization and modifying part of parameters. While TRADES exhibits better repair results on the convolutional neural network compared to the fully connected ones, its efficacy diminishes when addressing local robustness errors with a larger radius. Concerning $\mathrm{DD}$, AdvRepair and REASSURE outperform the other methods on all the models.
Table 2: Results of correcting violation of safety properties on ACAS Xu
Model RSR/RGR/FDD/Time/s
CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours
$N_{2,1}$ 59.2 100.0 100.0 100.0 100.0 60.1 97.5 100.0 100.0 100.0 39.1 50.3 64.2 9.2 0.0 28.1 4.3 572.3 18.3 25.3
$N_{2,2}$ 67.0 100.0 100.0 100.0 100.0 65.6 99.0 99.7 100.0 100.0 56.5 87.6 96.5 9.9 0.0 23.6 3.8 545.6 18.6 18.8
$N_{2,3}$ 100.0 100.0 100.0 100.0 100.0 99.9 99.6 100.0 100.0 100.0 11.1 92.0 80.1 9.3 0.0 14.2 2.9 522.8 17.8 18.2
$N_{2,4}$ 100.0 100.0 100.0 100.0 100.0 98.8 99.9 100.0 100.0 100.0 10.5 97.6 67.6 8.8 0.0 13.9 3.1 515.4 17.9 18.5
$N_{2,5}$ 98.4 100.0 100.0 100.0 100.0 97.8 98.9 64.1 100.0 100.0 9.6 90.9 91.9 5.4 0.0 28.8 3.7 522.0 16.8 19.0
$N_{2,6}$ 59.4 100.0 100.0 100.0 100.0 54.7 98.5 87.3 100.0 100.0 5.1 92.8 94.6 2.7 0.0 29.2 4.4 518.1 17.2 24.4
$N_{2,7}$ 100.0 100.0 100.0 100.0 100.0 100.0 99.3 92.5 100.0 100.0 5.9 97.0 88.0 1.7 0.0 11.7 5.0 520.6 16.1 30.4
$N_{2,8}$ 100.0 100.0 100.0 100.0 100.0 100.0 97.8 93.7 100.0 100.0 6.3 95.7 86.1 2.1 0.0 12.7 4.4 526.1 15.8 18.1
$N_{2,9}$ 97.8 100.0 100.0 100.0 100.0 98.1 90.5 92.6 100.0 100.0 13.5 94.9 88.9 1.5 0.0 24.4 5.1 512.2 15.7 24.2
$N_{3,1}$ 100.0 100.0 100.0 100.0 100.0 100.0 97.8 99.2 100.0 100.0 8.7 90.9 79.2 7.3 0.0 12.7 6.3 521.1 17.2 18.5
$N_{3,2}$ 100.0 100.0 100.0 100.0 100.0 100.0 99.3 95.3 100.0 100.0 8.9 12.2 48.6 8.6 0.0 12.2 3.0 510.9 16.7 18.3
$N_{3,3}$ 100.0 100.0 100.0 100.0 100.0 99.6 94.3 98.8 100.0 100.0 35.7 66.4 68.6 8.3 0.0 24.6 4.3 516.6 17.4 18.5
$N_{3,4}$ 100.0 100.0 100.0 100.0 100.0 100.0 98.1 97.7 100.0 100.0 18.9 97.5 96.0 9.6 0.0 23.3 3.1 513.9 15.6 23.5
$N_{3,5}$ 99.6 100.0 100.0 100.0 100.0 99.4 97.7 80.7 100.0 100.0 13.0 98.8 95.4 5.8 0.0 33.8 3.6 528.0 17.5 17.8
$N_{3,6}$ 100.0 100.0 100.0 100.0 100.0 99.9 98.7 97.2 100.0 100.0 9.8 94.3 96.6 2.7 0.0 16.1 4.6 639.5 16.8 17.5
$N_{3,7}$ 100.0 100.0 100.0 100.0 100.0 100.0 91.6 81.8 100.0 100.0 5.5 96.4 80.0 2.2 0.0 20.6 5.0 515.4 16.9 18.6
$N_{3,8}$ 100.0 100.0 100.0 100.0 100.0 100.0 98.8 98.3 100.0 100.0 5.0 97.5 88.8 1.6 0.0 15.2 4.1 579.9 17.9 24.8
$N_{3,9}$ 100.0 100.0 100.0 100.0 100.0 99.6 96.0 49.2 100.0 100.0 9.2 97.0 86.1 1.9 0.0 16.2 4.9 774.4 16.8 19.0
$N_{4,1}$ 100.0 100.0 100.0 100.0 100.0 99.9 98.3 98.8 100.0 100.0 40.1 96.2 93.2 7.8 0.0 14.5 3.5 549.4 17.6 24.5
$N_{4,3}$ 98.4 100.0 100.0 100.0 100.0 97.5 99.4 99.8 100.0 100.0 20.1 91.1 87.5 7.0 0.0 28.0 2.9 506.7 17.1 17.7
$N_{4,4}$ 100.0 100.0 100.0 100.0 100.0 100.0 98.7 99.9 100.0 100.0 4.3 82.0 85.2 8.4 0.0 9.5 3.1 535.8 16.9 17.9
$N_{4,5}$ 100.0 100.0 100.0 100.0 100.0 100.0 99.8 94.6 100.0 100.0 33.6 46.4 94.8 6.0 0.0 9.8 5.0 544.8 18.2 17.7
$N_{4,6}$ 100.0 100.0 100.0 100.0 100.0 100.0 98.2 97.9 100.0 100.0 11.2 89.6 86.0 2.6 0.0 19.9 5.4 548.2 15.9 24.1
$N_{4,7}$ 77.6 100.0 100.0 100.0 100.0 78.1 98.8 85.6 100.0 100.0 3.9 96.3 96.2 1.6 0.0 38.4 4.5 544.0 16.0 24.5
$N_{4,8}$ 99.8 100.0 100.0 100.0 100.0 99.7 97.4 77.1 100.0 100.0 4.7 97.9 88.2 1.6 0.0 36.0 4.9 542.7 14.8 30.2
$N_{4,9}$ 99.8 100.0 100.0 100.0 100.0 99.8 95.0 64.0 100.0 100.0 7.7 98.4 90.1 3.3 0.0 21.2 5.0 545.8 16.4 23.8
$N_{5,1}$ 100.0 100.0 100.0 100.0 100.0 100.0 90.5 97.6 100.0 100.0 14.0 98.4 96.9 6.9 0.0 19.8 3.5 541.3 17.1 18.6
$N_{5,2}$ 100.0 100.0 100.0 100.0 100.0 99.9 99.3 100.0 100.0 100.0 4.1 91.8 94.1 7.2 0.0 13.3 4.0 543.4 17.2 18.3
$N_{5,3}$ 100.0 100.0 100.0 100.0 100.0 99.9 88.6 99.7 100.0 100.0 8.8 88.3 94.8 6.0 0.0 17.4 3.5 542.9 16.8 18.7
$N_{5,4}$ 99.6 100.0 100.0 100.0 100.0 98.9 95.9 99.5 100.0 100.0 17.7 98.0 91.9 7.1 0.0 16.9 2.9 540.7 16.1 18.2
$N_{5,5}$ 100.0 100.0 100.0 100.0 100.0 100.0 97.0 99.7 100.0 100.0 5.4 94.8 94.4 3.9 0.0 8.6 3.4 518.9 15.9 18.3
$N_{5,6}$ 100.0 100.0 100.0 100.0 100.0 100.0 99.2 98.8 100.0 100.0 6.0 95.7 93.7 9.6 0.0 13.1 4.7 519.5 15.9 30.0
$N_{5,7}$ 97.2 100.0 100.0 100.0 100.0 97.7 99.1 98.6 100.0 100.0 4.9 95.3 87.7 1.8 0.0 21.2 5.0 517.3 16.3 24.2
$N_{5,8}$ 99.2 100.0 100.0 100.0 100.0 98.9 99.7 95.5 100.0 100.0 5.2 99.2 80.0 1.8 0.0 28.3 4.3 517.9 15.9 24.1
$N_{5,9}$ 100.0 100.0 100.0 100.0 100.0 100.0 96.9 91.7 100.0 100.0 5.4 97.2 27.0 2.0 0.0 22.6 4.8 515.8 15.4 24.3
Avg 95.8 100.0 100.0 100.0 100.0 95.5 97.3 92.2 100.0 100.0 13.4 88.8 85.1 5.2 0.0 20.0 4.2 540.9 16.8 21.4
The results of correcting safety property violations on ACAS Xu are presented in Table 2. For the task of correcting safety property violation, AdvRepair, PRDNN, REASSURE, and ART successfully construct a provable repair, resulting in a 100% $\mathrm{RSR}$. In contrast, CARE attains an average $\mathrm{RSR}$ of 95.8%. Regarding $\mathrm{RGR}$, all baseline methods exhibit notable performance, with ART and our approach showcasing the highest performance at 100%. In terms of $\mathrm{FDD}$, our approach achieves the most favorable performance with zero, surpassing CARE, ART, PRDNN, and REASSURE. The results highlight the effectiveness of our approach in achieving provable repairs and maintaining high fidelity.
Answer to RQ1: AdvRepair consistently outperforms the baselines in both local robustness repair and safety property violation correction tasks. It achieves a high repair success rate coupled with lower drawdown, showcasing superiority over the alternative approaches. The efficiency of AdvRepair is acceptable.
4.3 Generalization
In this experiment, we evaluate the generalization of the repaired network and its defense capabilities against new adversarial attacks, as shown in Table 3. AdvRepair outperforms the other baselines in terms of $\mathrm{RGR}$, showcasing superior performance. Notably, PRDNN, APRNN and TRADES also demonstrate some level of generalization, especially on the convolutional neural network, while the remaining tools exhibit relatively poor results. In terms of $\mathrm{DSR}$, AdvRepair exhibits superior performance compared to the other tools. TRADES is the only alternative showing a certain defensive capability against adversarial attacks on FNN_big and CNN, particularly when the radius is small. When evaluating $\mathrm{DGSR}$, we find that AdvRepair once again achieves the best performance. CARE, PRDNN, APRNN and TRADES demonstrate limited defense capabilities against FNN_big and CNN, especially when the radius is small. Notice that $\mathrm{DGSR}$ is the most significant metric for assessing generalization to global inputs. These results are sufficient to show that AdvRepair achieves a breakthrough in repairing adversarial attacks.
Answer to RQ2: AdvRepair exhibits robustness defense against new adversarial attacks, surpassing the defensive performance of other tools. AdvRepair exhibits significant advantages in terms of the generalization of repairing adversarial attacks.
Table 3: Results of generalization and defense against new adversarial attacks
Model $r$ $n$ RGR/DSR/DGSR/%
CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
FNN_ small 0.05 50 7.6 44.4 47.2 10.8 80.0 10.0 100.0 0.0 0.0 2.0 0.0 26.0 6.0 100.0 8.9 3.2 2.3 0.0 9.4 8.9 90.0
100 2.2 48.7 46.7 9.1 0.0 10.0 100.0 1.0 0.0 0.0 0.0 2.0 20.0 100.0 9.3 5.1 0.8 0.0 9.8 11.3 96.4
200 5.0 42.6 16.1 8.5 0.0 11.0 100.0 2.0 0.0 0.5 0.0 3.5 10.0 100.0 8.6 1.1 1.3 0.0 9.8 10.1 96.6
500 6.6 39.0 10.3 7.2 0.0 11.0 100.0 1.4 0.0 3.8 0.0 2.8 9.6 100.0 8.3 0.5 5.3 0.0 9.8 8.9 96.6
1000 12.7 41.3 10.9 2.0 11.0 100.0 2.3 0.0 0.4 3.9 14.4 100.0 6.1 0.2 0.9 0.0 9.8 11.3 96.6
0.1 50 0.8 24.2 32.0 2.4 20.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 8.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.3
100 1.7 29.8 21.9 1.8 10.0 9.0 100.0 0.0 0.0 0.0 0.0 13.0 9.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 95.0
200 1.1 32.7 12.9 1.4 5.0 12.0 100.0 0.0 0.0 0.0 0.0 10.5 11.5 100.0 0.0 0.0 0.0 0.0 9.8 10.1 95.2
500 1.7 37.5 17.3 1.4 8.0 14.0 100.0 0.0 0.0 0.0 0.0 10.0 10.0 100.0 0.0 0.0 0.1 0.0 9.8 9.8 95.3
1000 2.3 42.5 15.2 13.0 14.0 100.0 0.0 0.0 1.2 10.0 10.0 100.0 0.0 0.0 0.2 0.0 9.8 9.8 96.5
0.3 50 0.0 24.2 37.4 0.8 20.0 8.0 94.8 0.0 0.0 0.0 0.0 8.0 8.0 92.0 0.0 0.0 0.0 0.0 9.8 9.7 86.4
100 7.7 22.0 18.4 0.4 10.0 11.0 99.0 0.0 0.0 0.0 0.0 13.0 11.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.2
200 0.0 26.1 8.8 0.5 5.0 8.0 99.5 0.0 0.0 0.0 0.0 10.5 9.5 100.0 0.0 0.0 0.0 0.0 9.8 9.6 90.8
500 0.4 30.4 12.8 0.3 8.0 10.0 99.5 0.0 0.0 0.0 0.0 10.0 10.4 99.8 0.0 0.0 0.0 0.0 9.8 10.3 92.0
1000 0.6 31.9 12.5 13.0 11.0 99.7 0.0 0.0 0.0 9.8 9.8 99.7 0.0 0.0 0.0 0.0 9.8 9.8 96.6
FNN_ big 0.05 50 6.6 63.0 64.8 8.4 100.0 24.0 100.0 2.0 2.0 8.0 0.0 82.0 24.0 100.0 37.7 38.3 31.7 0.0 53.5 11.3 97.0
100 10.9 63.3 67.4 9.6 90.0 9.0 100.0 7.0 1.0 11.0 0.0 60.0 9.0 100.0 35.7 39.8 19.0 0.0 54.9 10.3 97.1
200 10.7 58.9 64.6 7.3 69.5 6.0 100.0 4.5 2.5 2.5 0.0 54.5 5.0 100.0 37.9 33.2 12.8 0.0 57.0 8.9 97.1
500 10.6 54.8 66.2 78.8 10.0 100.0 4.8 0.6 0.0 52.8 7.4 100.0 35.4 29.7 1.7 0.0 61.2 10.1 97.2
1000 11.8 54.6 68.4 79.2 20.0 100.0 7.1 0.4 0.0 56.1 19.2 100.0 36.8 25.5 1.8 0.0 66.2 11.3 97.2
0.1 50 20.2 52.6 44.2 3.4 0.0 8.0 100.0 2.0 0.0 0.0 0.0 0.0 8.0 100.0 2.6 0.7 1.1 0.0 1.8 9.7 97.1
100 2.0 47.5 46.3 3.2 20.0 11.0 100.0 0.0 0.0 0.0 0.0 7.0 11.0 100.0 1.8 0.9 0.0 0.0 4.0 9.8 97.1
200 19.2 49.6 46.8 2.5 22.5 10.0 100.0 1.5 0.0 0.0 0.0 6.0 10.5 100.0 2.6 0.8 0.1 0.0 4.9 9.8 97.1
500 1.2 50.6 49.3 23.6 15.0 100.0 0.0 0.2 0.0 2.8 9.0 100.0 1.2 0.2 0.0 0.0 5.4 8.9 97.2
1000 1.0 51.1 49.9 42.1 11.0 100.0 0.0 0.0 0.0 7.2 10.5 100.0 1.4 0.1 0.0 0.0 9.2 10.3 97.2
0.3 50 0.0 42.2 20.4 2.8 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.0
100 0.1 43.4 27.8 1.4 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.7
200 0.1 40.8 31.0 1.6 11.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.3 97.2
500 0.3 36.9 36.0 9.2 9.0 100.0 0.0 0.0 0.0 0.0 10.2 100.0 0.0 0.0 0.0 0.0 0.0 10.3 97.2
1000 0.4 36.5 37.1 15.0 12.0 100.0 0.0 0.0 0.0 0.0 11.8 99.9 0.0 0.0 0.0 0.0 0.0 11.3 97.2
CNN 0.05 50 27.0 69.6 67.6 47.2 100.0 0.0 100.0 0.0 0.0 12.0 0.0 100.0 18.0 100.0 70.6 37.5 70.4 0.0 87.0 9.8 96.6
100 29.4 74.5 65.6 45.7 100.0 60.0 100.0 1.0 0.0 10.0 0.0 86.0 27.0 100.0 70.6 29.4 69.2 0.0 88.2 9.7 98.3
200 27.6 79.8 62.5 50.9 97.0 0.0 100.0 1.0 0.0 3.0 0.0 84.0 20.5 100.0 70.6 10.4 70.6 0.0 89.9 9.7 98.3
500 30.8 89.0 62.3 92.4 100.0 1.4 0.0 0.8 87.0 100.0 71.4 0.0 91.2 98.3
1000 32.9 92.2 63.8 90.8 100.0 1.5 0.0 0.2 86.8 100.0 65.7 0.0 92.1 98.3
0.1 50 12.2 53.0 41.0 10.4 80.0 0.0 100.0 2.0 0.0 2.0 0.0 50.0 8.0 100.0 8.3 0.0 8.7 0.0 44.9 10.3 98.2
100 22.2 69.1 46.0 14.7 88.0 10.0 100.0 1.0 0.0 0.0 0.0 57.0 11.0 100.0 8.2 0.0 8.6 0.0 48.5 9.6 98.2
200 21.8 73.2 46.1 13.0 96.5 20.0 100.0 0.5 0.0 0.5 0.0 62.0 11.0 100.0 8.2 0.0 9.2 0.0 53.3 9.8 98.2
500 30.4 84.5 47.0 93.2 100.0 0.2 0.0 0.0 66.6 99.8 8.4 0.0 62.4 98.2
1000 24.6 90.3 48.7 94.7 100.0 0.6 0.0 0.0 70.1 99.9 4.6 0.0 67.8 98.3
0.3 50 0.0 44.8 70.0 4.0 0.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 12.0 100.0 0.0 0.0 0.0 0.0 0.0 9.8 98.3
100 0.0 59.9 75.5 2.7 13.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 7.0 100.0 0.0 0.0 0.0 0.0 0.0 8.9 98.3
200 0.0 69.4 78.2 3.6 12.5 0.0 100.0 0.0 0.0 0.0 0.0 0.0 7.0 100.0 0.0 0.0 0.0 0.0 0.0 8.9 98.3
500 1.3 79.6 73.3 12.6 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.3
1000 1.1 85.9 70.6 16.0 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.2
4.4 Scalability
We assess the scalability of AdvRepair on VGG19 and ResNet18 for local robustness repair using CIFAR-10 through the feature-space repair. The experimental results are presented in Table 4. Regarding $\mathrm{RSR}$, AdvRepair, PRDNN and APRNN achieve 100%. TRADES also performs commendably, although it slightly trails behind the provable tools. Notably, CARE exhibits an inability to repair the local robustness properties on the CIFAR-10 dataset. In terms of $\mathrm{RGR}$, AdvRepair demonstrates superior performance, with TRADES also achieving better results than other tools. As for $\mathrm{DD}$, AdvRepair outperforms the others except in scenarios where the number of buggy input is 1 000. In this situation, TRADES attains the best performance. This phenomenon is reasonable, since TRADES tends to perform better in the DNN accuracy when the number of samples in adversarial training are getting large. Looking at $\mathrm{DSR}$, we find that AdvRepair exhibits significantly better performance than the other tools, with the latter lacking the capability to defend against adversarial attacks on the original samples. For $\mathrm{DGSR}$, AdvRepair achieves the same level as the accuracy, surpassing the other tools. Although TRADES also performs comparatively well, it remains notably lower than AdvRepair. The running time of AdvRepair is acceptable, regarding the far superior repair performance and generalization.
Answer to RQ3: AdvRepair demonstrates good scalability in repairing local robustness on large-scale DNNs, outperforming the other tools.
Table 4: Results of repairing local robustness on CIFAR-10
VGG19 ResNet-18
$r=4/255$ $r=8/255$ $r=4/255$ $r=8/255$
50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000
RSR/CARE 8.0 5.0 4.0 11.0 10.9 2.0 4.0 2.0 2.2 2.4 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
PRDNN 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
TRADE 96.0 95.0 95.0 97.4 96.7 96.0 92.0 92.5 95.6 96.0 100.0 99.0 100.0 99.8 99.5 100.0 96.0 99.5 99.4 96.9
APRNN 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
Ours 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
RGR/CARE 6.6 4.2 3.6 11.7 11.8 2.0 3.6 2.7 2.4 2.9 2.6 2.7 9.1 12.2 6.7 3.8 2.6 2.3 2.7 2.3
PRDNN 69.8 68.2 68.5 58.0 66.6 61.4 68.0 68.6 67.2 59.8 57.8 54.9
TRADE 84.0 90.0 85.0 97.8 98.0 80.0 90.0 100.0 98.0 99.0 100.0 90.0 100.0 100.0 99.1 100.0 90.0 100.0 100.0 97.3
APRNN 69.2 68.9 70.9 75.5 80.4 63.2 65.2 68.2 66.1 70.7 70.6 65.1 62.0 62.7 68.2 65.0 55.6 53.0 49.6 59.0
Ours 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
DD/CARE 0.1 0.2 0.2 0.2 0.3 0.1 0.2 0.2 0.1 0.3 0.1 0.1 0.2 0.3 0.2 0.1 0.0 0.0 0.1 0.1
PRDNN 13.6 24.8 30.3 19.8 40.5 50.2 2.1 2.1 2.0 14.0 13.7 12.2
TRADE 46.0 35.3 29.8 14.7 -3.3 43.9 37.8 29.7 11.8 -2.7 43.0 51.7 33.0 12.9 -11.2 50.4 51.4 29.2 13.7 -8.9
APRNN 14.2 20.0 29.3 48.7 46.4 24.0 39.5 61.9 75.5 82.1 1.8 1.8 2.2 3.1 6.1 5.5 4.2 7.6 12.7 19.5
Ours 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DSR/CARE 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
PRDNN 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
APRNN 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TRADE 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Ours 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
DGSR/CARE 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
PRDNN 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
APRNN 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TRADE 3.6 3.7 2.9 3.4 3.4 4.1 3.7 3.2 2.3 2.6 5.7 6.7 6.0 5.4 5.2 5.6 5.0 5.4 5.8 5.2
Ours 83.6 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3
Time/s CARE 47.8 64.4 106.5 380.4 928.2 36.9 48.2 96.0 234.1 443.0 45.5 58.9 105.5 243.5 384.9 37.1 75.4 101.6 180.7 285.7
PRDNN 1731.8 2640.6 6307.4 1160.4 2210.2 5126.9 731.8 1794.9 3307.5 715.4 1300.2 5757.3
TRADE 68.7 114.8 194.6 415.4 814.6 67.2 111.1 187.2 417.6 806.8 57.1 107.8 186.7 405.0 794.3 57.9 107.8 184.1 408.1 790.6
APRNN 188.9 288.2 451.3 1645.6 6211.4 209.2 318.6 492.4 1750.4 7787.1 307.7 746.1 812.2 1338.1 1710.8 580.9 990.8 863.2 1556.0 1988.4
Ours 238.7 471.6 963.0 2516.9 5441.6 229.9 470.7 950.1 2485.0 5436.0 265.8 525.1 1050.3 2586.4 5262.3 264.7 523.1 1042.8 2627.8 5325.1
4.5 Monotonicity w.r.t. the number of buggy behaviors
We systematically evaluate the performance of the repair results concerning the number of buggy behaviors in the local robustness repair task. We consider five distinct configurations: 50, 100, 200, 500 and 1000. The comparative analysis encompasses the performance of various baselines alongside our approach across these configurations. AdvRepair not only showcases effectiveness in limiting the number of buggy behaviors but also excels when faced with a growing number of buggy behaviors. In contrast, APRNN encounters challenges in successfully repairing the FNN_small model when the number of buggy behaviors exceeds 100, and REASSURE experiences memory overflow issues when confronted with 1000 buggy behaviors.
Answer to RQ4: Even in scenarios involving a large number of buggy behaviors, AdvRepair demonstrates efficient and effective removal of buggy behaviors while preserving the original performance of the models. Comparative analysis indicates that AdvRepair leverages the available buggy points to achieve superior repair results compared to alternative approaches.
5 Related work
Provable DNN repair. Our approach falls within the category of provable DNN repair methods. The methods most closely related to ours are ART [39] and REASSURE [22]. In comparison to REASSURE, our approach fundamentally differs in its repair objectives. REASSURE focuses on fixing activation patterns of neural networks, while we directly guarantee the properties, whose input constraints may encompass a large number of activation patterns. In this way, our method exhibits superior generalization compared to REASSURE. Although ART also utilizes training in its repair process, it directly tunes parameters in the original neural network. In contrast, our approach involves training patch modules specifically for different properties, enhancing the specialization of our repairs and, consequently, their effectiveness. Other methods such as PRDNN [60] and APRNN [63] formulate the repair problem in linear programming. However, their effectiveness diminishes in properties with high-dimensional polytopes, such as the robustness of image classifications.
Heuristic DNN repair. Utilizing heuristic algorithms such as particle swarm optimization and differential evolution, CARE [61] and Arachne [59] aim to localize the neurons responsible for the fault. VeRe [41] focuses on providing formal verification guidance for fault localization and defining target intervals for repair synthesis. DeepRepair [82] and few-shot guided mix [51], for instance, expand the set of negative samples to generate additional training data. Tian et al. [64], on the other hand, augment the existing data by introducing real-world environmental effects, like fog, to the samples. To the best of our knowledge, DL2 [21] is the first paper to combine logical constraints and loss function, but it does not adopt convex approximation to give the loss function a certified guarantee.
DNN verification. In 2010, [50] proposed the first method for DNN verification. Over the past decade, various approaches of DNN verification has been proposed including techniques such as constraint solving [9, 25, 31, 38, 47, 33, 18, 28], abstract interpretation [24, 57, 80, 56, 36, 55], linear relaxation [30, 49, 79, 74, 7], global optimisation [52, 17, 53], CEGAR [2, 19, 48], reduction to two-player games [75, 77], and star-set abstraction [67, 66]. These method offer provable estimations of DNN robustness. Moreover, statistical approaches, presented in [72, 73, 5, 76, 6, 11, 27, 44, 12, 37], prove to be more efficient and scalable, particularly suited for intricate DNN structures, allowing for the establishment of quantifiable robustness at a specified confidence level. Certified training [4, 45, 46] uses convex approximation in training the loss function, but it is only used to enhance the robustness of DNN, without making it accessible to repair properties in the training procedure.
6 Conclusion
We introduce AdvRepair, a novel approach for addressing adversarial attacks using limited data. Our method provides provable and targeted repairs within the robustness neighborhood, enabling the generalization of this defense to other inputs. In terms of efficiency, scalability, and generalization, our approach surpasses existing methods. For future work, we plan to integrate BaB-based verification methods to enhance the performance of our approach, improving both precision and offering an alternative refinement approach. Additionally, our patch-based repair framework has the potential to evolve into a black-box repair method.
Appendix 0.A Proof of Theorem 3.4
Theorem 0.A.1
Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
$\mathcal{L}^{*}(\varphi):=\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0),$
where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
Proof
As $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is a linear function with respect to $x$, we can calculate the maximum of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on ${B(x_{i},r)}$ as follows:
$\max_{x\in B(x_{i},r)}\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$
$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\beta_{\ell}$
$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\cdot(x_{i}+r)+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\cdot(x_{i}-r)+\beta_{\ell}$
Then we have
$\mathcal{L}(\varphi)(x)=\sum_{\ell\neq\ell_{0}}\max(\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},0)=0,\forall x\in B(x_{i},r)$
$\iff$ $\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x-r\cdot\bm{1})+\beta_{\ell},0)=0$
$\iff$ $\mathcal{L}^{*}(\varphi)=0.$
Therefore, if $\mathcal{L}^{*}(\varphi)=0$, we have
$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}\leq 0,$ (2)
then the property $\varphi$ holds. ∎
Appendix 0.B Impact of Patch Module Size and Quantity on Efficiency
In this experiment, we assess the efficacy of size and quantity of patch modules in the context of repairing local robustness properties. We conduct a comparative analysis of AdvRepair’s performance across various sizes and quantities, presenting the experimental outcomes in Table 5. Two distinct patch module sizes, namely small and large, are considered. The small patch module consists of three hidden layers with 64 neurons in the first two layers and 16 neurons in the last layer, while the large patch module comprises three hidden layers with 256 neurons in the first two layers and 64 neurons in the last layer. Additionally, we explore two different numbers of patch modules: property-guided patch modules and untargeted label patch modules.
Table 5: The efficacy of patch module scale and quantity
Model $r$ $n$ RSR/RGR/DSR/DGSR/%
PS PL LS LL PS PL LS LL PS PL LS LL PS PL LS LL
FNN_ small 0.05 50 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 90.0 90.0 89.9 89.8
100 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.4 96.4 96.1 96.0
200 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
500 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.3 96.3 89.5 89.5
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.0 96.3 94.7 94.8
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.2 95.1 94.9 94.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.3 95.2 96.3 96.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.5 96.5 96.3 96.3
0.3 50 94.0 94.0 94.0 94.0 94.8 94.8 94.8 94.8 92.0 92.0 92.0 92.0 86.4 79.9 44.6 44.5
100 99.0 99.0 99.0 99.0 99.0 99.0 99.0 99.0 100.0 99.0 99.0 100.0 96.2 96.2 85.8 86.1
200 99.5 99.5 99.5 99.5 99.5 99.5 99.5 99.5 100.0 99.0 97.5 98.0 90.8 91.2 85.9 85.9
500 99.4 99.4 99.8 99.8 99.5 99.5 99.8 99.8 99.8 99.6 100.0 100.0 92.0 92.1 89.5 89.7
1000 99.6 99.6 99.9 99.9 99.7 99.7 100.0 100.0 99.7 99.6 100.0 100.0 96.6 96.5 94.2 94.2
FNN_ big 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.0 97.0 96.8 96.8
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.2 96.8 96.9
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.9 96.8
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 96.8 96.8
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.0
0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.0 96.1 89.1 89.1
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.7 96.8 89.1 89.1
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 89.3 89.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 95.9 95.9
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.9 100.0 100.0 97.2 97.2 97.1 97.1
CNN 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.7
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.0 98.0
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.1 98.0
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.5 99.5 98.2 98.2 98.0 98.1
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.8 99.8 99.8 99.8 98.2 98.2 98.0 98.1
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.8 99.8 99.8 98.3 98.3 98.0 98.0
0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.3 98.3
The performance of AdvRepair is evaluated across different combinations of patch module scales and quantities, including small property-guided patch module (PS), large property-guided patch modules (PL), small untargeted label repairing patch modules (LS), and large untargeted label repairing patch modules (LL). Notably, the experimental findings indicate that both sizes of the patch modules exhibit commendable performance, with minimal impact observed on the repair performance due to the network size. Concerning the quantity of patch modules, the property-guided patch modules outperform the untargeted label repairing patch modules in defending against adversarial attacks within $D_{t}$. We hypothesize that this superiority is attributed to the specialization inherent in property-guided repair.
Appendix 0.C Experimental Details
Table 6: Network architectures and accuracies of DNN trained on MNIST dataset
Name Accuracy/Model structure
FNN_small 96.6 linear layer of 50 hidden units
linear layer of 50 hidden units
linear layer of 50 hidden units
linear layer of 50 hidden units
linear layer of 32 hidden units
linear layer of 10 hidden units
FNN_big 98.3 linear layer of 200 hidden units
linear layer of 200 hidden units
linear layer of 200 hidden units
linear layer of 200 hidden units
linear layer of 32 hidden units
linear layer of 10 hidden units
CNN 97.2 Conv2d(1, 16, 4, stride=2, padding=1)
linear layer of 100 hidden units
linear layer of 10 hidden units
The network architectures and accuracies of the DNN trained on MNIST are detailed in Table 6. For CIFAR-10, the accuracies of VGG19 and ResNet18 are 93.4% and 88.3%, respectively. In Algorithm 1, the maximum number of iterations $M$ is set to 25. In Algorithm 2, the maximum number of epochs $R$, learning rate $\eta$, and selection number $K$ are set to 10, 10, and 800, respectively. The attack methods in AutoAttack have a maximum of 10 iterations, and the number of restarts is set to 1.

### v2

@@
-AdvRepair: Provable Repair of Adversarial Attack
+Patch Synthesis for Property Repair of Deep Neural Networks
 Abstract
-Deep neural networks (DNNs) are increasingly deployed in safety-critical domains, but their vulnerability to adversarial attacks poses serious safety risks.
-Existing neuron-level methods using limited data lack efficacy in fixing adversaries due to the inherent complexity of adversarial attack mechanisms, while adversarial training, leveraging a large number of adversarial samples to enhance robustness, lacks provability.
-In this paper, we propose AdvRepair, a novel approach for provable repair of adversarial attacks using limited data.
-By utilizing formal verification, AdvRepair constructs patch modules that, when integrated with the original network, deliver provable and specialized repairs within the robustness neighborhood.
-Additionally, our approach incorporates a heuristic mechanism for assigning patch modules, allowing this defense against adversarial attacks to generalize to other inputs.
-AdvRepair demonstrates superior efficiency, scalability and repair success rate.
-Different from existing DNN repair methods, our repair can generalize to general inputs, thereby improving the robustness of the neural network globally, which indicates a significant breakthrough in the generalization capability of AdvRepair.
-1 Introduction
-In the last few years, deep neural networks (DNNs) have made significant advancements in various domains, including computer vision [3], natural language processing [16], and speech recognition [71].
-However, the adoption of DNNs in safety-critical domains has been slow due to concerns regarding their dependability.
-One major concern is the vulnerability of DNNs to adversarial attacks [43, 14], where adversaries can manipulate input data in a way that is imperceptible to humans but can cause the model to make incorrect decisions.
+Deep neural networks (DNNs) are prone to various dependability issues, such as adversarial attacks, which hinder their adoption in safety-critical domains.
+Recently, NN repair techniques have been proposed to address these issues while preserving original performance by locating and modifying guilty neurons and their parameters.
+However, existing repair approaches are often limited to specific data sets and do not provide theoretical guarantees for the effectiveness of the repairs.
+To address these limitations, we introduce PatchPro, a novel patch-based approach for property-level repair of DNNs, focusing on local robustness.
+The key idea behind PatchPro is to construct patch modules that, when integrated with the original network, provide specialized repairs for all samples within the robustness neighborhood while maintaining the network’s original performance.
+Our method incorporates formal verification and a heuristic mechanism for allocating patch modules, enabling it to defend against adversarial attacks and generalize to other inputs.
+PatchPro demonstrates superior efficiency, scalability, and repair success rates compared to existing DNN repair methods, i.e., realizing provable property-level repair for 100% cases across multiple high-dimensional datasets.
+I Introduction
+In recent years, deep neural networks (DNNs) have achieved significant advancements in various domains, including computer vision [1], natural language processing [2], and speech recognition [3].
+Despite these advancements, the adoption of DNNs in safety-critical domains has been slow due to concerns regarding their dependability.
+A major concern is the vulnerability of DNNs to adversarial attacks [4, 5], where adversaries can manipulate input data in ways that are imperceptible to humans but can cause the model to make incorrect decisions.
 This vulnerability poses serious safety risks in applications such as autonomous vehicles [6] and medical diagnosis [7].
-Therefore, it is crucial to develop and deploy DNN models with safety and reliability requirements in mind to fully realize their potential in safety-critical domains.
-The repair of DNNs involves rectifying errors in the network by modifying its architecture or parameters.
-These errors can arise from various sources, including training errors [10], adversarial attacks [26, 13], backdoor attacks [40], and distribution drift [78, 62].
-Regular and efficient repair during DNN deployment is crucial to mitigate the risks associated with these errors and ensure the performance and reliability of the network in practical applications.
-Traditional machine learning approaches for DNN repair include techniques such as adversarial training [24, 23, 65], input sanitization, fine-tuning, transfer learning [15, 81], and data augmentation [82, 51, 42].
-These methods rely on a large amount of data and may exhibit suboptimal performance in situations with limited or low-quality data.
-Neuron-level fault localization and repair methods [61] aim to improve the accuracy and reliability of DNN repair by identifying and correcting errors at the individual neuron level.
-By leveraging these neuron-level techniques, researchers and practitioners can achieve more precise and reliable DNN repair with limited data.
-At present, the efficacy of neuron-level repair methods utilizing only a limited number of samples for robust defense against adversarial attacks remains limited.
-These methods demonstrate effectiveness primarily in the context of backdoor attacks or individual safety properties.
-This limitation arises from the inherent complexity of adversarial attack mechanisms, which often exhibit a higher degree of intricacy compared to backdoor attacks.
-Moreover, diverse adversarial attacks targeting distinct samples correspond to a unique dense mixture of features [1], rendering it arduous to generalize parameter adjustments at the neuron level based on a small sample set.
-Furthermore, while techniques such as adversarial training leverage a large number of adversarial samples to bolster the robustness of neural networks, their efficacy lacks provability.
-Consequently, neural networks repaired using such techniques continue to harbor latent vulnerabilities in terms of robustness.
-Our proposed methodology aims to address the issue of adversarial attacks by leveraging a limited set of adversarial samples.
-Specifically, our approach focuses on rectifying these erroneous samples while simultaneously maximizing accuracy and ensuring local robustness on these samples.
-Moreover, we aim to achieve this local robustness in a manner that generalizes across the entire dataset.
-To achieve these objectives, our method integrates an external indicator, a hard-coded characteristic function, which facilitates the allocation of patch modules in conjunction with the original network.
-This indicator identifies inputs within the local neighborhood of a specific sample and assigns the appropriate patch modules for repair.
-Additionally, for inputs outside the local neighborhood, a heuristic allocation mechanism is employed to enable generalization across the entire dataset.
-Each patch module constitutes a fully connected neural network structure.
-The core concept of our approach lies in the utilization of reachability analysis through a linear approximation for the provable training of patch modules.
-Specifically, we employ the verification method DeepPoly [58] to establish a linear approximation of the output neurons.
-This approximation is leveraged to determine the distance between the target behavior and the current behavior, serving as the loss function.
+Therefore, it is crucial to regularly update the DNN to mitigate the risks associated with these errors and ensure the reliability of the network in practical applications.
+DNN repair techniques [8, 9, 10] have been proposed to address the problem involving rectifying errors in the network by modifying its architecture or parameters.
+Compared to traditional methods such as adversarial training [11, 12, 13], input sanitization, fine-tuning, transfer learning [14, 15], and data augmentation [16, 17, 18], neuron-level fault localization and repair methods [8] offer a more targeted approach by identifying and correcting errors at the individual neuron level while not affecting the overall network performance.
+Despite significant advancements in this field, several limitations remain.
+First, neuron-level repair techniques, which rely on a limited number of samples, often struggle to provide robust defense against adversarial attacks due to the inherent complexity of these attacks compared to simpler threats like backdoor attacks.
+Adversarial attacks involve intricate mixtures of features [19], making it difficult to generalize parameter adjustments from a small dataset.
+Second, existing repair methods typically focus on specific data for repair and fail to generalize to the property level (e.g., local robustness), which limits their effectiveness in addressing a broad range of adversarial scenarios.
+This motivates us to explore property repair of DNNs, specifically to fix the safety properties that neural networks violate in certain input regions.
+Third, while provable repair methods such as PRDNN [20], REASSURE [21], and APRNN [22] can conduct property-based error correction, they achieve this by ensuring that outputs meet constraints on the vertices of the input region as a polyhedron.
+However, the number of vertices increases exponentially with data dimensionality, which makes these methods inefficient for high-dimensional data.
+In this work, we aim to address these limitations by proposing a novel patch-based method to achieve provable repair on the property level.
+Specifically, we focus on correcting potential adversarial samples in the infinite set of high-dimensional points within a certain neighborhood of these error samples, i.e., satisfying local robustness property.
+To achieve this, our key idea is to use formal verification to help construct a separate patch module (in the form of a fully connected neural network structure) for each neighborhood outside of the original network.
+Such a patch-based method allows us not only to ensure that the infinite set of points within the error sample’s neighborhood repaired but also to maintain the performance of the original network unaffected.
+To construct such a patch, we utilize reachability analysis through linear relaxation for the provable training of these patch modules.
+Specifically, we use the verification method DeepPoly [23] to create a linear relaxation of the output neurons.
+This approximation is employed to determine the distance between the targeted behavior and the current behavior, serving as the loss function.
 The patch modules are then trained to minimize this distance.
-Once the loss function reaches zero, the patch modules provide provable repairs for adversarial attacks within the perturbation region.
-By employing an external indicator to identify faulty inputs, adversarial attacks within the same perturbation region can be repaired using the same patch module.
-Inputs outside the perturbation range of limited adversarial samples are assigned by the indicator, leveraging a heuristic allocation mechanism, thereby ensuring the generalization of the repair.
-Through the comprehensive framework of our method, we effectively enhance the resilience of the network against adversarial attacks while preserving accuracy.
-By leveraging reachability analysis and patch modules, we provide provable repairs for adversarial samples, thereby improving the robustness of the network.
-The contributions of this paper can be summarized as follows:
-  • We propose a novel approach, named AdvRepair, specifically designed for provable repair of adversarial attacks using limited data.
-This work represents a significant breakthrough in achieving provable repair for adversarial robustness.
-Our method introduces patch modules that are seamlessly integrated into the original network architecture.
-By combining a loss function derived from formal methods, we train these patch modules to provide provable repairs for local robustness.
-Through a heuristic allocation mechanism, our approach successfully achieves the generalization of local robustness repair.
-  • To address the efficiency challenges associated with formal verification in large-scale DNNs, we leverage patch modules to perform repairs at the feature space of the networks.
-This strategy enables our method to scale effectively across various network architectures, ensuring efficient and practical implementation.
-  • We conduct a comprehensive evaluation of AdvRepair on three diverse datasets and multiple DNN architectures.
-Through extensive comparisons with state-of-the-art repair and adversarial training approaches, our method consistently demonstrates superior efficiency and scalability.
-Moreover, it exhibits remarkable generalization capabilities to handle general inputs, thereby significantly enhancing the overall robustness of the network.
-This advancement in generalization capabilities represents a substantial contribution of our method.
-Overall, our proposed approach, AdvRepair, offers a pioneering solution for provable repair of adversarial attacks using limited data.
-It not only addresses the challenges posed by formal verification in large-scale DNNs but also achieves remarkable generalization capabilities, leading to improved network robustness.
-2 Preliminary
+Once the loss function reaches zero, the patch modules will offer provable repairs for adversarial attacks within the perturbation region.
+To ensure that each patch module addresses specific neighborhoods, our approach employs an external indicator that identifies inputs within a particular sample’s local neighborhood and assigns the appropriate patch modules for repair.
+This indicator allows the same patch module to effectively repair adversarial attacks within the same perturbation region.
+For adversarial samples outside the repaired property, the indicator utilizes a heuristic allocation mechanism to assign suitable patch modules, thereby enhancing the generalization of the repair.
+This comprehensive framework not only improves the network’s resilience against adversarial attacks but also maintains its accuracy.
+By leveraging reachability analysis and dedicated patch modules, we provide provable repairs for adversarial samples, significantly boosting the network’s robustness.
+Additionally, to extend this local robustness across the entire dataset, we integrate the external indicator with the original network, ensuring that all samples, including those beyond the initially repaired property, benefit from heuristic patch module allocation.
+This approach effectively combines local and global repair strategies to enhance overall network robustness.
+We summarize our contributions as follows:
+  • We introduce PatchPro, a novel approach that integrates a loss function derived from formal verification to train patch modules within the original network.
+This method effectively repairs high-dimensional infinite point sets within the polyhedron neighborhood of adversarial samples while preserving the network’s performance.
+The approach also includes a heuristic allocation mechanism, which ensures the generalization of local robustness repair by seamlessly integrating patch modules into the original network architecture.
+  • To tackle the efficiency challenges of formal verification in large-scale DNNs, we utilize patch modules to perform repairs in the feature space of the networks.
+This strategy enables our method to scale effectively across various network architectures, ensuring both efficient and practical implementation.
+  • We thoroughly evaluate PatchPro on three diverse datasets and multiple DNN architectures.
+Through extensive comparisons with state-of-the-art repair and adversarial training techniques, our method consistently demonstrates superior efficiency, scalability, and generalization capabilities.
+It shows significant improvement in handling general inputs, thereby greatly enhancing the overall robustness of the network.
+II Preliminary
 In this section we recall some basic notations of DNN repair.
@@
 In many applications, DNNs are serving for classification tasks.
-In such a classification DNN ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$, every output dimension corresponds to a classification label, and the one with the maximum output value is the classification results that the DNN $N$ gives, i.e., $C_{N}(x)=\arg\max_{1\leq i\leq n_{L}}N(x)_{i}$, where $N(x)_{i}$ is the $i$th entry of the vector $N(x)$.
+In such a classification DNN ${N}\colon\mathbb{R}^{n_{0}}\rightarrow\mathbb{R}^{n_{L}}$, every output dimension corresponds to a classification label, and the one with the maximum output value is the classification result that the DNN $N$ gives, i.e., $C_{N}(x)=\arg\max_{1\leq i\leq n_{L}}N(x)_{i}$, where $N(x)_{i}$ is the $i$th entry of the vector $N(x)$.
 The notion of safety properties pertains to assertions that guarantee the absence of undesirable behavior.
 Within the context of DNNs, a safety property demands that a DNN operates correctly within a specified input range.
-The local robustness of a DNN refers to its ability to maintain stable and consistent predictions in the vicinity of its training data points, even in the presence of small perturbations or variations in the input.
+Definition 1.
+A safety property is a triple $({N},{X},{Q})$, where $N$ is a DNN, ${X}\subseteq\mathbb{R}^{d_{\text{in }}}$ and ${Q}\subseteq\mathbb{R}^{d_{\text{out }}}$ are the subset of input and output spaces of the neural network ${N}$.
+The property $({N},{X},{Q})$ is satisfied if and only if ${N}(x)\in{Q}$ for all $x\in{X}$.
+A local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, where a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius.
+Formally, it can be defined as follows:
+The local robustness of a DNN refers to its ability to maintain stable and consistent predictions in the vicinity of its in-distribution data points, even in the presence of small perturbations or variations in the input.
 Here a neighborhood of an input $x_{0}$ is usually defined as a closed ball $B(x_{0},r):=\{x\in\mathbb{R}^{n_{0}}\mid\|x-x_{0}\|_{\infty}\leq r\}$, where $\|\cdot\|_{\infty}$ is the $L_{\infty}$-norm, and $r>0$ is the radius.
 Formally, a local robustness property of a classification DNN $N$ requires that for any input $x$ in a given neighborhood $B(x_{0},r)$ of an input $x_{0}$, its classification should always be consistent with $x_{0}$, i.e., $\forall x\in B(x_{0},r),C_{N}(x)=C_{N}(x_{0})$.
-We denote this local robustness property as $(N,B(x_{0},r))$; it is obviously a safety property.
-An input $x^{*}\in B(x_{0},r)$ is an adversarial sample of the DNN $N$, if $C_{N}(x^{*})\neq C_{N}(x_{0})$, i.e., $x^{*}$ is a counterexample which violates the local robustness property $(N,B(x_{0},r))$.
+We denote this local robustness property as $(N,B(x_{0},r))$, i.e.
+$(N,B(x_{0},r))=(N,B(x_{0},r),\{y\in\mathbb{R}^{n_{L}}\mid y_{i}<y_{C_{N}(x_{0})},i=1,2,\dots,n_{L}\})$, and thus it is a safety property.
 In this work, we focus on the problem of repairing adversarial attacks with limited adversarial samples.
@@
 Now we formally state the problem of repairing adversarial attacks as follows:
-Given a DNN $N$ and a set of adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$, where $n$ can be significantly smaller than the size of the dataset, and each $x_{i}^{*}$ is obtained by an adversarial attack on an input $x_{i}$ with a given radius $r$, we need to construct a DNN $F$ which is locally robust on every $B(x_{i},r)$ while the accuracy is maintained and local robustness of other inputs is potentially improved.
-3 Methodology
+Given a DNN $N$ and a set of adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$, where $n$ can be significantly smaller than the size of the training set, and each $x_{i}^{*}$ is obtained by an adversarial attack on an input $x_{i}$ with a given radius $r$, we need to construct a DNN $F$ which is locally robust on every $B(x_{i},r)$ while the accuracy is maintained and local robustness of other inputs is potentially improved.
+III Methodology
 We focus on fixing adversarial attacks with limited data in this work, and the aims of the repair are at least threefold: For a given radius $r>0$,
@@
   • the accuracy of the DNN is maintained, and for as many input $x$ in the dataset as possible, the DNN is locally robust in $B(x,r)$.
-Here the first objective is the fundamental requirement showing that the repair is successful on the given adversarial samples $\{x_{i}^{*}\}_{i=1}^{n}$.
-The second objective aims to ensure the local generalizability of the repair, whereby the absence of adversarial examples is guaranteed within the vicinity of the generated adversarial examples, while the third objective focuses on the global generalizability of the repair, wherein it simultaneously improves the robustness against all potential inputs in the dataset without significantly compromising the accuracy.
-The current neural network repair methods under limited data conditions primarily involve parameter adjustments at the neuron level.
-While these methods demonstrate high efficacy against straightforward backdoor attacks, they face significant challenges in effectively addressing complex adversarial attacks.
-Even if the erroneous behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ are successfully repaired, such repairs lack the ability to generalize either locally or globally.
-Inspired by Reassure [22], we propose a patch-based repair method.
+In this work, we propose a patch-based repair method.
 A patch refers to a specific modification or alteration made to a software system or codebase; it is a discrete set of changes applied to fix a bug, enhance functionality, or address security vulnerabilities.
@@
 Furthermore, for other inputs in the dataset, we can heuristically match patches that maximize their robustness, thereby endowing this repair approach with global generalization.
-In this section, we propose a patch-based repair method named AdvRepair to defend from adversarial attacks with limited data.
-3.1 Structure of the repaired DNN
-The main produce of the repaired network is depicted in Fig.
+In this section, we propose a patch-based repair method named PatchPro to defend from adversarial attacks with limited data.
+III-A Structure of the repaired DNN
+The main produce of the repaired network is shown in Fig.
 1.
-For an input $x\in\mathbb{R}^{n_{0}}$, the role of the indicator is to select an appropriate patch that, when applied, yields the sum of the output of the patch module and the output of the original DNN for $x$.
+For an input $x\in\mathbb{R}^{n_{0}}$, the role of the indicator is to select the appropriate patches that, when applied, yields the sum of the outputs of the patch modules and the output of the original DNN for $x$.
 This sum represents the output obtained after the repair process.
-Figure 1: Framework of AdvRepair
-The definition of an indicator function is standard:
-Definition 1
-Let $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties. The indicator function $\mathcal{I}_{\mathcal{C}}\colon\mathbb{R}^{n_{0}}\to\{0,1\}^{m}$ outputs in the $j$th entry as
-$\mathcal{I}_{\mathcal{C}}(x)_{j}=\left\{\begin{array}[]{ll}1,&\text{ if }x\models X_{j},\\ 0,&\text{ otherwise},\end{array}\right.$
-where $j=1,\ldots,m$.
+The indicator function is defined as:
+Figure 1: The architecture of a DNN repaired by PatchPro.
+It contains multiple patch networks for each input properties.
+Each of them is enabled or disabled according to the allocation signal “1” or “0” determined by the indicator.
+The blue lines highlights the indicator’s workflow.
+The final output is the sum of the outputs of all the enabled patches and the original network.
+Definition 2.
+Let $\mathcal{C}=(X_{1},\ldots,X_{m})$ be a finite sequence of input properties.
+The indicator function $\mathcal{I}_{\mathcal{C}}\colon\mathbb{R}^{n_{0}}\to\{0,1\}^{m}$ outputs in the $j$th entry, where $j\in\{1,\dots,m\}$, as
+$\mathcal{I}_{\mathcal{C}}(x)_{j}=\left\{\begin{array}[]{ll}1,&\text{ if }x\models X_{j},\\ 0,&\text{ otherwise}.\end{array}\right.$
+Typically an input property $X_{i}$ is a subset of $\mathbb{R}^{n_{0}}$, and we define $x\models X_{i}$ iff $x\in X_{i}$.
 Upon classification by the indicator, a set of patch modules is deployed to perform the repair.
@@
 The implementation of a patch module is a fully connected neural network in this work.
-For an input $x$ that is not within the input set of any indicator, i.e., $x\notin\bigcup_{i=1}^{n}B(x_{i},r)$, it cannot be identified by the indicators, but there is still potential adversarial attack in $B(x,r)$.
-In this situation, we heuristically allocate a patch module to this input to defend from adversarial attacks.
+For an input $x$ that does not satisfy any input properties, i.e., $x\notin\bigcup_{i=1}^{n}B(x_{i},r)$, there is no existing specific patch module, but it may still suffer from adversarial attacks within its neighborhood.
+In this situation, we heuristically allocate some patch modules to this input to defend from adversarial attacks.
 Formally, the structure of the repaired DNN is as follows:
-Definition 2
+Definition 3.
 A repaired DNN is a tuple $F=(N,\mathcal{C},\mathcal{P},\tau)$, where
@@
   • $\mathcal{P}=(P_{1},\ldots,P_{m})^{\mathrm{T}}$ is a finite sequence of patch modules, each of which is a fully connected neural network,
-  • and $\tau\colon\mathbb{R}^{n_{0}}\setminus\bigcup_{i=1}^{n}B(x_{i},r)\to 2^{\{1,\ldots,m\}}$ is a patch allocation function.
-The semantics of the repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$ is a function
+  • and $\tau\colon\mathbb{R}^{n_{0}}\setminus\bigcup_{i=1}^{n}X_{i}\to 2^{\{1,\ldots,m\}}$ is a patch allocation function.
+The semantics of the repaired DNN $F$ is a function:
 $F\colon\mathbb{R}^{n_{0}}$ $\to\mathbb{R}^{n_{L}},$
-$x$$\mapsto\left\{\begin{array}[]{ll}\mathcal{I}_{\mathcal{C}}(x)^{\mathrm{T}}\mathcal{P}(x)+N(x),&\text{ if }x\in\bigcup_{i=1}^{n}B(x_{i},r),\\ \sum_{j\in\tau(x)}P_{j}(x)+N(x),&\text{ otherwise},\end{array}\right.$
+$x$$\mapsto\left\{\begin{array}[]{ll}N(x)+\mathcal{I}_{\mathcal{C}}(x)^{\mathrm{T}}\mathcal{P}(x),&\text{ if }x\in\bigcup_{i=1}^{n}X_{i},\\ N(x)+\sum_{j\in\tau(x)}P_{j}(x),&\text{ otherwise},\end{array}\right.$
 where $\mathcal{P}(x)=(P_{1}(x),\ldots,P_{m}(x))^{\mathrm{T}}$.
-Generally there are three ways to determine the number of the patch modules and define the indicator:
-  • Property-guided.
-The property-guided way sets $\mathcal{C}_{\mathrm{P}}=\{B(x_{i},r)\mid i=1,\ldots,n\}$, i.e., the indicator judges which robustness region the input belongs to, and the corresponding patch module defends against adversarial attacks in this specific region.
-  • Untargeted label.
-A simple label-guided way takes every classification label as an input property.
-i.e., $\mathcal{C}_{\mathrm{L}}=\{1,\ldots,n_{L}\}$, where $x\models\ell$ iff $C_{N}(x)=\ell$.
-Assuming that $x$ is correctly classified by $N$, the patch module to defend against adversarial attacks in $B(x,r)$ is determined by $C_{N}(x)$, and it works for fixing potential adversarial attacks of all the robustness regions where the ground truth is $C_{N}(x)$.
-  • Targeted label.
-We consider the set of label pairs $\mathcal{C}_{\mathrm{L}}=\{(\ell_{0},\ell_{1})\in\{1,\ldots,n_{L}\}^{2}\mid\ell_{0}\neq\ell_{1}\}$, where $x\models(\ell_{0},\ell_{1})$ iff $N(x)_{\ell_{0}}>N(x)_{\ell_{1}}>N(x)_{\ell}$ for all $\ell\neq\ell_{0},\ell_{1}$, i.e., the indicator determines which label $\ell_{1}$ is the most likely when attacking the ground truth $\ell_{0}$, assuming that $x$ is correctly classified by $N$, and the corresponding patch module fixes adversarial attacks of this specific pattern.
-The property-guided way has advantages when the number $n$ of the given adversarial samples is small, while the label-guided ways work well in the situation when the number $n_{L}$ of the output labels is small.
-For the label-guided ways, the patch allocation function $\tau$ can be simply defined as the corresponding indicator, since this indicator can naturally generalize to other input.
-In the following, we present our approach in the property-guided way, and it can be easily adapted to the label-guided ways.
-3.2 Training the patch modules
-The first challenge is to train the patch modules, which are each implemented as a fully connected neural network, to repair each robustness property $(N,B(x_{i},r))$ with a provable guarantee.
+In this work, we set $\mathcal{C}=\{B(x_{i},r)\mid i=1,\ldots,n\}$, i.e., the indicator judges which robustness region the input belongs to, and the corresponding patch module defends against adversarial attacks in this specific region.
+III-B Training the patch modules
+The first challenge is to train the patch modules, which are each implemented as a fully connected neural network, to repair each robustness property $(F,B(x_{i},r))$ with a provable guarantee.
 To achieve provability, we employ formal verification in training these patch modules.
@@
 This idea has been widely adopted in DNN repair [24, 25].
-In this work, we employ DeepPoly [58] as the verification engine.
+In this work, we employ DeepPoly [23] as the verification engine [^1] Also there are several other verification tools based on abstract interpretation, such as CROWN [26] and DeepZ [27], which vary in precision, affecting loss function evaluations.
+We choose DeepPoly since it provides a good balance between efficiency and precision..
 DeepPoly uses abstract interpretation to give every neuron an upper/lower bound in the form of an affine function, where only variables in previous layers occur, and the numerical bound can be derived by propagating backwards these affine upper/lower bounds to the input layer.
-Recall that a local robustness property $(N,B(x_{i},r))$ holds iff for any $x\in B(x_{i},r)$ and $\ell\neq\ell_{0}=C_{N}(x_{i})$, $N(x)_{\ell}<N(x)_{\ell_{0}}$, so DeepPoly calculates the abstraction of $N$ on $B(x_{i},r)$ for every neuron and the expressions $N(x)_{\ell}-N(x)_{\ell_{0}}$ for $\ell\neq\ell_{0}$.
-From the abstraction, we can obtain an affine function in which there are only input variables as a sound upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$, i.e.,
-$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},$ (1)
+Recall that a local robustness property $(F,B(x_{i},r))$ holds iff for any $x\in B(x_{i},r)$ and $\ell\neq\ell_{0}=C_{F}(x_{i})$, $F(x)_{\ell}<F(x)_{\ell_{0}}$, so DeepPoly calculates the abstraction of $F$ on $B(x_{i},r)$ for every neuron and the expressions $F(x)_{\ell}-F(x)_{\ell_{0}}$ for $\ell\neq\ell_{0}$.
+From the abstraction, we can obtain an affine function in which there are only input variables as a sound upper bound of $F(x)_{\ell}-F(x)_{\ell_{0}}$ on $B(x_{i},r)$, i.e.,
+$\forall x\in B(x_{i},r),F(x)_{\ell}-F(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},$ (1)
 where $\alpha_{\ell}\in\mathbb{R}^{n_{0}}$ and $\beta_{\ell}\in\mathbb{R}$ are constants.
 It is easy to obtain the numerical upper bound of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on the box region $B(x_{i},r)$.
-If this upper bound is negative for every $\ell\neq\ell_{0}$, then the local robustness property $(N,B(x_{i},r))$ is verified to be true by DeepPoly.
+If this upper bound is negative for every $\ell\neq\ell_{0}$, then the local robustness property $(F,B(x_{i},r))$ is verified to be true by DeepPoly.
 Therefore, it is natural to use this upper bound as the loss function to train the corresponding patch module.
-Definition 3
-For a local robustness property $\varphi=(N,B(x_{i},r))$, we define the safety violated loss function of $\varphi$ as
+Definition 4.
+For a local robustness property $\varphi=(F,B(x_{i},r))$, we define the safety violated loss function as
 $\mathcal{L}(\varphi)(x)=\sum_{\ell\neq\ell_{0}}\max(\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},0),$
-where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is the upper bound of $N(x)_{\ell}-N(x)_{\ell_{0}}$ on $B(x_{i},r)$ given by DeepPoly, as shown in Eq.
+where $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is the upper bound of $F(x)_{\ell}-F(x)_{\ell_{0}}$ on $B(x_{i},r)$ given by DeepPoly, as shown in Eq.
 (1).
 For local robustness properties $\varphi_{1},\ldots,\varphi_{k}$ which share the same ground truth label $\ell_{0}$, we define $\mathcal{L}(\bigwedge_{i=1}^{k}\varphi_{i})=\sum_{i=1}^{k}\mathcal{L}(\varphi_{i})$.
-For a local robustness property $\varphi=(N,B(x_{i},r))$, its safety violated loss function $\mathcal{L}(\varphi)$ being 00 implying that $\varphi$ is verified to be true by DeepPoly.
-Theorem 3.1
-Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
-$\mathcal{L}^{*}(\varphi):=\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
-where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
-When we train a patch module $P$ on the DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$, the maximum of the safety violated loss function of a local robustness property $\varphi$, namely $\mathcal{L}^{*}(\varphi)$, is parameterized with the weights $w$ in $P$, and this $\mathcal{L}^{*}(\varphi)(w)$ is the loss function for training this patch $P$.
-By employing the standard gradient descent method, we can train the patch $P$, until the loss reaches 00 or the number of epochs reaches a threshold $R$.
-We claim that, even if $\mathcal{L}^{*}(\varphi)(w)$ does not reach 00, the repair does not necessarily fail, since the upper bound $\mathcal{L}^{*}(\varphi)$ given by DeepPoly is not exact; as it decreases in the training process, $F$ is growing more likely to satisfy $\varphi$.
-Also, to improve the precision of $\mathcal{L}^{*}(\varphi)$ obtained from DeepPoly, we employ the input interval partitioning technique from ART [39].
+For a local robustness property $\varphi=(F,B(x_{i},r))$, its safety violated loss function $\mathcal{L}(\varphi)$ being 00 implies that $\varphi$ is verified to be true by DeepPoly.
+Theorem 1.
+Let $\varphi=(F,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
+$\mathcal{L}^{*}(\varphi):=$ $\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot$
+$(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
+where $\mathrm{elmax}$ and $\mathrm{elmin}$ are the element-wise $\max$ and $\min$ operation, $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
+The notion that $\mathcal{L}^{*}(\varphi)$ is the expansion of $\mathcal{L}(\varphi)(x)$ after taking corresponding values on the boundary of the ball.
+To improve the precision of $\mathcal{L}^{*}(\varphi)$ obtained from DeepPoly, we employ the input interval partitioning technique from ART [25].
 It selects the partition dimension by computing the multiplication of the partial derivative of the safety violated loss function $\mathcal{L}(\varphi)(x)$ and the size of the input interval in the corresponding dimension, and bisects the box region over the dimension with the maximum score.
-Algorithm 1 AdvRepair (Property-guided)
-1:
-2:Original DNN $N$, pairs of input and its adversarial examples $\{(x_{i},x_{i}^{*})\}_{i=1}^{n}$, and radius $r>0$
-3:
-4:A repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$
-5:$\mathcal{C}\leftarrow(B(x_{1},r),\ldots,B(x_{n},r))$, $\mathrm{iter}\leftarrow 0$
-6:$\mathrm{Init}(\mathcal{P})$triangleright Initialize the patch modules $\mathcal{P}=(P_{1},\ldots,P_{n})$
-7:$D[\cdot,\cdot]\leftarrow\{\}$triangleright A dictionary where $D[i]$ stores properties fixed with $P_{i}$
-8:for $i\leftarrow 1~{}\text{to}~{}n$ do
-9:     $D[i]\leftarrow\{(F,B(x_{i},r))\}$
-10:$E\leftarrow\{1,\ldots,m\}$triangleright Record the properties not fixed yet
-11:while $\mathrm{iter}<M$ do $\triangleright$ $M$: Maximum number of iterations
-12:     $\mathrm{iter}\leftarrow\mathrm{iter}+1$
-13:     for $j\in E$ do
-14:         $(P_{j},T,\mathrm{Repaired})\leftarrow\textsc{Train}(N,P_{j},D[j])$triangleright Alg. 2
-15:         if $\mathrm{Repaired}=\mathbf{True}$ then
-16:              $E\leftarrow E\setminus\{j\}$
-17:         else
-18:              for $\varphi\in T$ do $\triangleright$ We write $\varphi=(F,X)$
-19:                  for $d\leftarrow 1~{}\text{to}~{}n_{0}$ do
-20:                       $\operatorname{score}_{d}\leftarrow\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|$
-21:                  $d^{*}\leftarrow\arg\max_{d}\operatorname{score}_{d}$
-22:                  $X_{1},X_{2}\leftarrow\mathrm{Bisect}(X,d^{*})$triangleright Bisection on the $d^{*}$th dimension
-23:                  $\varphi_{1}\leftarrow(F,X_{1})$, $\varphi_{2}\leftarrow(F,X_{2})$, $D[j]\leftarrow(D[j]\cup\{\varphi_{1},\varphi_{2}\})\setminus\{\varphi\}$
-24:     if $E=\{\}$ then
-25:         return $F$triangleright Repair with provable guarantee
-26:return $F$triangleright Repair without provable guarantee
+After partitioning, the property $\varphi$ is split into two new properties, whose input sets are two sub-boxes of the original input region, and the union of these two sub-boxes is the input set of $\varphi$.
+Algorithm 1 PatchPro
+1:Original DNN $N$, pairs of input and its adversarial examples $\{(x_{i},x_{i}^{*})\}_{i=1}^{n}$, and radius $r>0$
+2:A repaired DNN $F=(N,\mathcal{C},\mathcal{P},\tau)$
+3:$\mathcal{C}\leftarrow(B(x_{1},r),\ldots,B(x_{n},r))$, $\mathrm{iter}\leftarrow 0$
+4:$\mathrm{Init}(\mathcal{P})$triangleright Initialize the patch modules $\mathcal{P}=(P_{1},\ldots,P_{n})$
+5:$D[\cdot,\cdot]\leftarrow\{\}$triangleright A dictionary where $D[i]$ stores properties fixed with $P_{i}$
+6:for $i\leftarrow 1~{}\text{to}~{}n$ do
+7:    $D[i]\leftarrow\{(F,B(x_{i},r))\}$
+8:$E\leftarrow\{1,\ldots,n\}$triangleright Record the properties not fixed yet
+9:while $\mathrm{iter}<M$ do $\triangleright$ $M$: Maximum number of iterations
+10:    $\mathrm{iter}\leftarrow\mathrm{iter}+1$
+11:    for $j\in E$ do
+12:        $(P_{j},T,\mathrm{Repaired})\leftarrow\textsc{Train}(N,P_{j},D[j])$triangleright Alg. 2
+13:        if $\mathrm{Repaired}=\mathbf{True}$ then
+14:           $E\leftarrow E\setminus\{j\}$
+15:        else
+16:           for $\varphi\in T$ do $\triangleright$ We write $\varphi=(F,X)$
+17:               for $d\leftarrow 1~{}\text{to}~{}n_{0}$ do
+18:                  $\operatorname{score}_{d}\leftarrow\partial_{d}\mathcal{L}(\varphi)(x)\cdot\sup_{x,x^{\prime}\in X}|x_{d}-x_{d}^{\prime}|$
+19:               $d^{*}\leftarrow\arg\max_{d}\operatorname{score}_{d}$
+20:               $X_{1},X_{2}\leftarrow\mathrm{Bisect}(X,d^{*})$triangleright Bisection on the $d^{*}$th dimension
+21:               $\varphi_{1}\leftarrow(F,X_{1})$, $\varphi_{2}\leftarrow(F,X_{2})$, $D[j]\leftarrow(D[j]\cup\{\varphi_{1},\varphi_{2}\})\setminus\{\varphi\}$
+22:    if $E=\{\}$ then
+23:        return $F$triangleright Repair with provable guarantee
+24:return $F$triangleright Repair without provable guarantee
 Algorithm 2 Patch training
-1:
-2:Original DNN $N$, DNN $P$ as a patch module, and a finite set $U$ of local robustness properties
-3:
-4:The optimized patch module $P$, the set $T\subseteq U$ of properties to be refined, and whether the properties in $U$ have all been fixed with a provable guarantee
-5:function Train($N,P,U$)
-6:     $\mathrm{epoch}\leftarrow 0$
-7:     while $\mathrm{epoch}<R$  do $\triangleright$ $R$: maximum number of epochs
-8:         $\mathrm{epoch}\leftarrow\mathrm{epoch}+1$
-9:         $w\leftarrow w-\eta\cdot\nabla\mathcal{L}^{*}(\bigwedge U)(w)$triangleright $w$: the weights in $P$
-10:         if $\mathcal{L}^{*}(\bigwedge U)(w)=0$ then $\triangleright$ $\mathcal{L}^{*}(\bigwedge U):=\mathcal{L}^{*}(\bigwedge_{\varphi\in U}\varphi)=\sum_{\varphi\in U}\mathcal{L}^{*}(\varphi)$
-11:              return $(P,\{\},\mathbf{True})$
-12:     $T\leftarrow\mathrm{Slice}_{K}(\mathrm{ArgSort}_{\varphi}\{\mathcal{L}^{*}(\varphi)(w)\mid\mathcal{L}^{*}(\varphi)(w)>0,\varphi\in U\})$
-13:     return $(P,T,\mathbf{False})$
-The main algorithm of AdvRepair is shown in Alg.
+1:Original DNN $N$, DNN $P$ as a patch module, and a finite set $U$ of local robustness properties
+2:The optimized patch module $P$, the set $T\subseteq U$ of properties to be refined, and whether the properties in $U$ have all been fixed with a provable guarantee
+3:function Train($N,P,U$)
+4:    $\mathrm{epoch}\leftarrow 0$
+5:    while $\mathrm{epoch}<R$  do $\triangleright$ $R$: maximum number of epochs
+6:        $\mathrm{epoch}\leftarrow\mathrm{epoch}+1$
+7:        $w\leftarrow w-\eta\cdot\nabla\mathcal{L}^{*}(\bigwedge U)(w)$triangleright $w$: the weights in $P$
+8:        if $\mathcal{L}^{*}(\bigwedge U)(w)=0$ then $\triangleright$ $\mathcal{L}^{*}(\bigwedge U):=\sum_{\varphi\in U}\mathcal{L}^{*}(\varphi)$
+9:           return $(P,\{\},\mathbf{True})$
+10:    $T\leftarrow\mathrm{Slice}_{K}(\mathrm{ArgSort}_{\varphi}\{\mathcal{L}^{*}(\varphi)(w)\mid\mathcal{L}^{*}(\varphi)(w)>0,\varphi\in U\})$
+11:    return $(P,T,\mathbf{False})$
+The main algorithm of PatchPro is shown in Alg.
 1.
@@
 The repair process has at most $M$ iterations.
-In each iteration, we train the patches which does not yet provide provable repair, i.e., those whose index is in $E$.
+In each iteration, we train the patches which do not yet provide provable repair, i.e., those whose index is in $E$.
 The algorithm of training a specific patch module $P$ for repairing a set $U$ of properties is shown in Alg.
@@
 1).
-Although we are focusing on fixing adversarial attacks in this work, AdvRepair also works for repair safety properties of DNNs by making minor changes in Alg.
+Although we are focusing on fixing adversarial attacks in this work, PatchPro also works for repair safety properties of DNNs by making minor changes in Alg.
 1.
-In AdvRepair, we follow a classical way of defining the loss function with the linear relaxation obtained from DeepPoly, and incorporate it into our patch-based repair framework.
+In PatchPro, we follow a classical way of defining the loss function with the linear relaxation obtained from DeepPoly, and incorporate it into our patch-based repair framework.
 This combination has several advantages.
 First, every patch is responsible for repairing properties in a specific pattern, which makes it much easier to obtain a provable repair; even if the repair is not provable, the safety violated loss is significantly declining in the training process, and it is highly possible that the repaired DNN is locally robust.
-Different from the neuron-level repair, there is no modification on the original DNN $N$ in AdvRepair, and due to the design of the indicator, the patch modules does not affect the behaviors of other input before $\tau$ allocates a patch to it.
+Different from the neuron-level repair, there is no modification on the original DNN $N$ in PatchPro, and due to the design of the indicator, the patch modules do not affect the behaviors of other input before $\tau$ allocates a patch to it.
 This mechanism avoids the drawdown of accuracy, which is quite severe in many neuron-level repair methods.
@@
 In the following, we will present how to construct the patch allocation function $\tau$ to improve generalization.
-3.3 Patch allocation
-For the property-guided method in AdvRepair, we need to define a patch allocation function $\tau$ because the construction of patches in this way does not naturally generalize to other inputs.
+For a network $N$ with $L$ layers and a maximum of $n$ neurons per layer, the time complexity of running Deeppoly for one neuron is $O(n^{2}\cdot L)$ [23].
+The time complexity of running Algorithm 2 once, which involves backpropagation for gradient computation and update weights of the patch networks in line 5 of this algorithm, is $O(R\cdot n_{\mathrm{patch}}\cdot L_{\mathrm{patch}})$, where $n_{\mathrm{patch}}$ is the maximum number of neurons per layer in all patch networks, and $L_{\mathrm{patch}}$ is the number of layers in the patch networks.
+Furthermore, during the execution of PatchPro, difficult-to-repair properties are split into two easier-to-repair properties, which may increase the number of properties to repair up to a predefined upper limit, denoted as $K$.
+In summary, the overall time complexity of PatchPro is $O(R\cdot M\cdot K\cdot(n^{3}\cdot L^{2}+n_{\mathrm{patch}}^{3}\cdot L_{\mathrm{patch}}^{2}))$.
+III-C Patch allocation
+For generalization purposes, it is necessary to introduce a patch allocation function $\tau$, given that the current patch construction approach lacks inherent adaptability to in-distribution data points.
 This patch allocation function $\tau$ is the key to achieving good generalization of the patch modules trained locally to global inputs.
-To match the best patch modules for an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we need to extract some specific patterns of the patch modules which generalize to other inputs and their adversarial robustness.
-Such patterns may include:
-  • the ground truth of $x^{\prime}$, which can be represented with $\ell_{0}=C_{N}(x^{\prime})$,
-  • the targeted label, i.e., the index of the second largest entry in $N(x^{\prime})$,
-[…]
-In the implementation of $\tau$, the patches with the same ground truth label as $x^{\prime}$ is alternative for $x^{\prime}$, and we allocate the ten with the maximum cosine similarity between $N(x^{\prime})$ and $N(x_{i})$ to $x^{\prime}$.
-Here we do not match with the targeted label because it is possible that more than one labels can potentially attack $N$ in the robustness region $B(x^{\prime},r)$, and matching by the cosine similarity can cover this possibility.
+To match the best patch modules for an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we propose utilizing the prediction $\ell_{0}=C_{N}(x^{\prime})$ of the original network on $x^{\prime}$ as a guiding principle, with the aim of selecting patches that share the same ground truth label as $\ell_{0}$.
+Accordingly, we formally define the set of patch modules $\tau(x^{\prime})$ associated with input $x^{\prime}$ as $\tau(x^{\prime})=\{i\mid C_{N}(x^{\prime})=C_{N}(x_{i})\}$.
+This definition establishes that the set $\tau(x^{\prime})$ encompasses all indices $i$ satisfying the condition $C_{N}(x^{\prime})=C_{N}(x_{i})$, i.e., those instances where the prediction $\ell_{0}$ made by the original network on $x^{\prime}$ is identical to the prediction made on sample $x_{i}$ at index $i$.
+By determining the patch module set $\tau(x^{\prime})$ in this manner, we ensure consistency between the selected modules and the input $x^{\prime}$ in terms of their predictions by the original network.
+This, in turn, is expected to enhance the robustness repair effectiveness for the specific input $x^{\prime}$.
+Once a set of patch modules has been allocated to an input $x^{\prime}$, these modules effectively repair the entire robustness region of that input.
+Namely, we establish a new local robustness property such that any subsequent inputs $x\in B(x^{\prime},r)$ will also utilize the same set of patch modules as $x^{\prime}$, with the system’s response given by $F(x)=\sum_{j\in\tau(x^{\prime})}P_{j}(x)+N(x)$.
+In this context, the defense mechanism is established prior to the occurrence of adversarial attacks, obviating the need to re-allocate patch modules for each individual input $x$ within the robustness region, which would otherwise require computing $\tau(x)$ separately.
+This proactive construction of the defense ensures a consistent and efficient protection strategy against potential adversarial threats across the entire neighborhood of $x^{\prime}$, reinforcing the overall robustness.
 We seemingly have problem when $x^{\prime}$ is not correctly classified by $N$, because in this case we may allocate inappropriate patches.
@@
 Otherwise, the wrong classification of $x^{\prime}$ may result from backdoor attacks, biased training data, overfitting, etc, and such situations are beyond the scope of fixing adversarial attacks in this work.
-Once a set of patch modules has been allocated to an input, it works in the whole robustness region of this input.
-For an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we allocate it with $\tau(x^{\prime})$, a set of patch modules.
-After that, all the inputs $x\in B(x^{\prime},r)$ share the same set of patch modules as $x^{\prime}$, i.e., $F(x)=\sum_{j\in\tau(x^{\prime})}P_{j}(x)+N(x)$, instead of allocating $x$ with $\tau(x)$.
-In this sense, our defense is constructed before adversarial attacks.
-3.4 Repair in a feature space
-To repair an extremely large DNN with high input dimensionality, the memory and computational overheads associated with the linear relaxations obtained from formal verification are notably significant and, at times, prohibitive.
-Also, faced with a huge DNN, the precision of abstraction-based verification methods like DeepPoly is severely harmed during a large number of layer-by-layer propagation.
-Consequently, the key to adopting AdvRepair to large DNNs is to reduce the input dimensionality and the size of the neural network in formal verification.
+III-D Repair in a feature space
+Repairing large DNNs with high-dimensional inputs poses significant challenges due to the substantial memory and computational costs associated with formal verification techniques.
+These costs can often become prohibitive, especially when dealing with deep networks where the precision of abstraction-based methods, such as DeepPoly, deteriorates significantly over multiple layers of propagation.
+Consequently, the key to adopting PatchPro to large DNNs is to reduce the input dimensionality and the size of the neural network in formal verification.
 In the popular architectures of convolutional neural networks, the convolutional layers play the role of extracting important features from data, followed by several fully connected layers for classification according to these extracted features, so we call such a fully connected layer a feature space.
-Definition 4
+Definition 5.
 Let $N\colon\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{L}}$ be a DNN where the function from the $i$th layer to the $(i+1)$th is $f_{i}$, i.e., $N=f_{L-1}\circ\cdots\circ f_{0}$.
@@
 The dimensionality of a feature space is usually far smaller than that of the input layer.
-If we start with a feature space in employing DeepPoly, it does not need to calculate the abstraction of a great deal of neurons in the convolutional layers, thus getting its efficiency and precision enhanced.
-To conduct repair in a given feature space with AdvRepair, we first give an approximation of the feature space on every $B(x_{i},r)$.
+If we start with a feature space in employing DeepPoly, it does not need to calculate the abstraction of a great deal of convolutional layers, thus getting its efficiency and precision enhanced.
+Selecting the input feature space for repairing remains an open challenge, with limited theoretical research available.
+Our choice of the second-to-last fully connected layer is heuristic, based on the common understanding that earlier layers in a neural network are primarily responsible for feature extraction, while later layers handle decision-making.
+Repairing in the decision layer is more efficient and incurs lower abstraction costs.
+To conduct repair in a given feature space with PatchPro, we first give an approximation of the feature space on every $B(x_{i},r)$.
 Here we do not require that this approximation should be a sound abstraction of the real semantics of the feature space on $B(x_{i},r)$, because calculating such an over-approximation with high precision is quite time-consuming.
-Instead, we simply sample in $B(x_{i},r)$ and obtain the tightest box region that contains the behaviors of the samples in the feature space as the approximation.
+Instead, we simply sample in $B(x_{i},r)$ and obtain the tightest box region that contains the buggy behaviors of the samples in the feature space as the approximation.
+To identify samples in the feature space that are “close” to adversarial examples, we use the Projected Gradient Descent (PGD) method and Fast Gradient Sign Method (FGSM) to find samples in $B(x_{i},r)$, and obtain their corresponding feature space behaviors.
+Formally, the abstraction $(X_{(i)})_{j}$ can be defined as follows.
+Definition 6.
+Let $S$ be the set of samples obtained by applying PGD and FGSM to $N$ in $B(x_{i},r)$.
+For each sample $s\in S$, we compute its feature space behavior $N_{0,l}(s)$ at the $l$-th layer.
+The abstraction of $S$ in the feature space is defined as:
+$(X_{(i)})_{j}=[\min_{s\in S}(N_{0,l}(s))_{j},\max_{s\in S}(N_{0,l}(s))_{j}]$
+where $(X_{(i)})_{j}$ is the range of $j$-th dimension of $X_{(i)}$, $(N_{0,l}(s))_{j}$ is the $j$-th element of the $N_{0,l}(s)$, and $j=1,\ldots,n_{l}$.
+These samples, being actual adversarial examples in real-world scenarios, provide direct guidance for feature space repair.
 This approximation is not sound yet, but it does not mean that the inputs in $B(x_{i},r)$ whose behaviors in the feature space are not within this approximation will not be repaired by the corresponding patch module $P_{i}$, because the indicator $\mathcal{I}_{\mathcal{C}}$ still remains the same.
@@
 The correspondence of the robustness regions $B(x_{i},r)$ and the patch $P_{i}$ is always preserved, so $P_{i}$ is always working for repairing on $B(x_{i},r)$ with the correct classification label $\ell_{0}$.
-After we obtain an approximation $X_{(i)}$ for each $B(x_{i},r)$ in the feature space, we employ Alg.
+After obtaining an approximation $B_{S}$ for each $B(x_{i},r)$ in the feature space, we employ Alg.
 1, where the initialization of $D[i]$ in Line 5 is $(N_{l,L}+P_{i},X_{(i)})$ instead.
@@
 1, so that the loss $\mathcal{L}^{*}$ decreases as much as possible and we can achieve a better repair performance.
-Since we are here considering the feature space in AdvRepair, it is natural to take it into consideration the cosine similarity and the distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space.
-If the ground truth label of $x^{\prime}$ and $x_{i}$ matches, a high cosine similarity and a small distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space strongly indicate that the patch module $P_{i}$ for $x_{i}$ is effective in fixing adversarial attacks in $B(x^{\prime},r)$.
-In this case, the patches with the same ground truth label as $x^{\prime}$ are still alternative for $x^{\prime}$, but we allocate the ten patches with the maximum value of
-$\frac{\langle N_{0,l}(x^{\prime}),N_{0,l}(x_{i})\rangle}{\|N_{0,l}(x^{\prime})\|_{2}\cdot\|N_{0,l}(x_{i})\|_{2}}\cdot\|N_{0,l}(x^{\prime})-N_{0,l}(x_{i})\|_{2}.$
-Repair in a feature space is an effective way to make AdvRepair scale on large DNNs.
-Although we sacrifice provability, its repair performance and generalization are extremely ideal in practical scenarios, which we will see in our experimental evaluation.
-4 Experimental Evaluation
-In this section, we evaluate AdvRepair by answering the following research questions:
-RQ1:What is the overall performance of AdvRepair in repairing local robustness and correcting safety property violations?
-RQ2: Does the repaired DNN exhibit the capability to defend against new adversarial attacks?
-RQ3: Does AdvRepair have scalability to repair large networks?
-RQ4: What impact does the number of adversarial samples needed for repair have on the effectiveness of the repair process?
-4.1 Setup
-All the experiments are conducted on a machine equipped with an AMD EPYC 7763 64-Core Processor, 256 GB of memory, and an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory.
-Each experiment has a timeout set to 10 000 seconds.
-[^1] AdvRepair is available at an anonymous link: https://drive.google.com/drive/folders/1NU170UZ0XmYKgg6diKZc-XMcqsUAVEib?usp=drive˙link
-Dataset.
-We conduct evaluations of our approach on three datasets: MNIST [35], CIFAR-10 [34] and ACAS Xu [69, 29].
-On MNIST, we train two fully connected neural networks and one convolutional neural network.
-On CIFAR-10, we train a VGG19 [54] and a ResNet18 [20].
-We assess our approach across 35 ACAS Xu DNNs, which comprises six hidden layers, each housing 50 neurons.
-As documented in [32], these DNNs are expected to satisfy Property-2 but exhibit violations.
-Detailed model structures and hyperparameters are in Appendix 0.C.
-For the local robustness repair task, we subject the DNNs trained on MNIST and CIFAR-10 to PGD [43], generating 50, 100, 200, 500 and 1 000 adversarial samples.
-For MNIST, the radius $r$ is set to be 0.05, 0.1 and 0.3, while for CIFAR-10 it is set to be $\frac{4}{255}$ and $\frac{8}{255}$.
-On ACAS Xu, our objective is to repair the violation of Property-2 while preserving the original performance of the ACAS Xu DNNs.
-We sample 500 counterexamples as the buggy inputs for repair.
-Metrics.
-To assess generalization, we sample ten adversarial examples on each $B(x_{i},r)$ different from $x_{i}^{*}$ to form a generalization set $D_{\mathrm{g}}$ for MNIST and CIFAR-10, and use an independent set of 5 000 counterexamples for ACAS Xu.
-Besides these, we have an independent test set $D_{\mathrm{t}}$ of size 10 000, 10 000 and 5 000 for MNIST, CIFAR-10 and ACAS Xu, respectively.
-We employ AutoAttack [14] to attack the repaired DNN $F$ on $\{x_{i}\}_{i=1}^{n}$ and $D_{\mathrm{t}}$ with the same radius $r$.
-The metrics we use include repair success rate ($\mathrm{RSR}$), repair generalization rate ($\mathrm{RGR}$), drawdown ($\mathrm{DD}$), defense success rate ($\mathrm{DSR}$), and defense generalization success rate ($\mathrm{DGSR}$), defined as follows:
-$\mathrm{RSR}$ =|$=\frac{|\{\,i\mid C_{F}(x_{i}^{*})=C_{N}(x_{i})\,\}|}{n},\qquad\mathrm{RGR}=\frac{|\{{\,x\in D_{\mathrm{g}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{g}}\right|},$
-$\mathrm{DD}$ =|$=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{N}(x)=\ell_{x}\,\}|-|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{t}}\right|},$
-$\mathrm{DSR}$ =|$=\frac{|\{\,i\mid\exists x\in\mathrm{AA}(F,B(x_{i},r)),C_{F}(x)\neq C_{N}(x_{i})\,\}|}{n},$
-$\mathrm{DGSR}$ =|$=\frac{|\{\,x\in D_{\mathrm{t}}\mid\exists x^{\prime}\in\mathrm{AA}(F,B(x,r)),C_{F}(x^{\prime})\neq C_{N}(x)\,\}|}{\left|D_{\mathrm{t}}\right|},$
-where $\ell_{x}$ denotes the ground truth of $x$, and $\mathrm{AA}(\varphi)$ represents the set of potential adversarial samples obtained by attacking the local robustness property $\varphi$ with AutoAttack.
-On ACAS Xu, the metric Drawdown is replaced with Fidelity Drawdown ($\mathrm{FDD}$) because there is no labelled ground truth for inputs on ACAS Xu:
-$\mathrm{FDD}=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=C_{N}(x)\,\}|}{|D_{\mathrm{t}}|}.$
-The metrics $\mathrm{RSR}$ and $\mathrm{DD}$ (or $\mathrm{FDD}$) evaluates the overall performance of DNN repair by measuring the percentage of buggy inputs successfully repaired and how much the accuracy decreases, while $\mathrm{RGR}$, $\mathrm{DSR}$ and $\mathrm{DGSR}$ reflects how the repair generalizes to the robustness regions $B(x_{i},r)$ and to other inputs.
-Baselines.
-We compare AdvRepair with the state-of-the-art repair methods including CARE [61], APRNN [63], PRDNN [60], REASSURE [22] and ART [39].
-Since APRNN and PRDNN require selecting a layer to repair on, we traverse all eligible layers across each baseline, and report the layer that exhibit the best performance.
-Additionally, for the local robustness repair task, we compare our approach with an adversarial training method TRADES [83], where we select the trained model with the best performance in 200 epochs with their default parameters.
-Table 1: Results of repairing local robustness, where REASS and TRADE denote REASSURE and TRADES, respectively, and – means timeout or memory overflow.
+In performing feature layer repairs on large-scale networks, we persist in utilizing the $\tau$ function guided by the original network output to allocate corresponding repair modules.
+While the approximation $X_{(i)}$ is indeed effective in capturing commonality in buggy behaviors, establishing a direct association between the input neighborhood $B(x^{\prime},r)$ of a newly encountered point $x^{\prime}$ and its distribution within the feature layers of the extensive network proves challenging.
+Consequently, assigning repair modules based on the input neighborhood of $x^{\prime}$ emerges as a more targeted and reliable approach.
+Repair in a feature space is an effective way to make PatchPro scale on large DNNs.
+Although we sacrifice provability, its repair performance and generalization capabilities demonstrate remarkable effectiveness in practical scenarios, which we will see in our experimental evaluation.
+TABLE I: Results of repairing local robustness. We represent each tool’s name with their first three letters, such as CAR for CARE, APR for APRNN, etc. And “–” means timeout or memory overflow.
 Model $r$ $n$ RSR/DD/Time/s
-CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
+CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours
 FNN_ small 0.05 50 8.0 100.0 100.0 100.0 54.0 10.0 100.0 1.0 30.9 19.7 0.0 3.3 86.8 0.0 2.0 35.9 3.5 381.3 5.2 257.7 14.3
@@
 1000 0.6 100.0 0.0 10.0 14.0 100.0 0.1 42.4 73.0 86.8 86.8 0.0 8.0 3844.8 33.9 104.5 1733.2 768.1
-0.3 50 0.0 100.0 100.0 100.0 8.0 8.0 94.0 0.0 77.0 42.9 0.0 86.8 86.8 0.0 1.8 55.3 3.6 328.3 13.3 255.7 18.6
-100 6.0 100.0 100.0 100.0 13.0 11.0 99.0 7.0 74.1 61.5 0.0 86.8 87.0 0.0 2.6 115.7 9.5 670.3 13.4 322.3 52.0
-200 0.0 100.0 0.0 100.0 10.5 8.0 99.5 0.0 62.8 54.2 0.0 86.8 87.0 0.0 3.2 364.9 132.1 1306.2 23.2 405.6 115.8
-500 0.0 100.0 0.0 100.0 10.0 10.0 99.4 -0.1 57.1 65.7 0.0 86.8 86.3 0.0 4.8 1934.4 147.9 3248.1 52.5 555.8 255.9
-1000 0.0 100.0 0.0 9.8 11.0 99.6 0.0 58.3 83.9 86.8 86.8 0.0 7.4 3759.6 44.7 103.2 2116.1 935.8
+0.3 50 0.0 100.0 100.0 100.0 8.0 8.0 100.0 0.0 77.0 42.9 0.0 86.8 86.8 0.0 1.8 55.3 3.6 328.3 13.3 255.7 18.6
+100 6.0 100.0 100.0 100.0 13.0 11.0 100.0 7.0 74.1 61.5 0.0 86.8 87.0 0.0 2.6 115.7 9.5 670.3 13.4 322.3 52.0
+200 0.0 100.0 0.0 100.0 10.5 8.0 100.0 0.0 62.8 54.2 0.0 86.8 87.0 0.0 3.2 364.9 132.1 1306.2 23.2 405.6 115.8
+500 0.0 100.0 0.0 100.0 10.0 10.0 100.0 -0.1 57.1 65.7 0.0 86.8 86.3 0.0 4.8 1934.4 147.9 3248.1 52.5 555.8 255.9
+1000 0.0 100.0 0.0 9.8 11.0 100.0 0.0 58.3 83.9 86.8 86.8 0.0 7.4 3759.6 44.7 103.2 2116.1 935.8
 FNN_ big 0.05 50 6.0 100.0 100.0 100.0 98.0 24.0 100.0 2.0 0.7 3.6 0.0 0.4 85.8 0.0 1.7 89.5 11.2 1807.6 5.7 436.5 14.5
@@
 1000 0.0 100.0 100.0 0.6 100.0 0.1 56.5 2.2 0.4 0.0 16.8 423.8 1667.0 80.4 1282.6
-4.2 Repair performance
-For the overall repair performance, we focus on $\mathrm{RSR}$, $\mathrm{DD}$ and running time, which indicate whether the buggy behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ can be fixed, whether the accuracy is maintained, and the efficiency, respectively.
-The results of the local robustness repair task on MNIST are summarized in Table 1.
-In the experiments of repairing local robustness, PRDNN, REASSURE and AdvRepair all achieve a perfect 100% $\mathrm{RSR}$, attributed to their provable constructions.
-APRNN, while provable, exhibits limitations in effectively repairing specific instances within the FNN_small model.
-Despite being a provable method, ART is less effective in repairing tasks involving high-dimensional inputs.
-CARE and TRADES, which do not have provable guarantee, demonstrate lower $\mathrm{RSR}$ compared to the provable approaches.
-CARE exhibits suboptimal repair performance for all the three models, possibly due to the intricate relationship between local robustness repair and the numerous parameters in the DNNs.
-The challenge lies in resolving such errors through fault localization and modifying part of parameters.
-While TRADES exhibits better repair results on the convolutional neural network compared to the fully connected ones, its efficacy diminishes when addressing local robustness errors with a larger radius.
-Concerning $\mathrm{DD}$, AdvRepair and REASSURE outperform the other methods on all the models.
-Table 2: Results of correcting violation of safety properties on ACAS Xu
+TABLE II: Results of correcting violation of safety properties on ACAS Xu
 Model RSR/RGR/FDD/Time/s
-CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours
+CAR PRD REA ART Ours CAR PRD REA ART Ours CAR PRD REA ART Ours CAR PRD REA ART Ours
 $N_{2,1}$ 59.2 100.0 100.0 100.0 100.0 60.1 97.5 100.0 100.0 100.0 39.1 50.3 64.2 9.2 0.0 28.1 4.3 572.3 18.3 25.3
@@
 Avg 95.8 100.0 100.0 100.0 100.0 95.5 97.3 92.2 100.0 100.0 13.4 88.8 85.1 5.2 0.0 20.0 4.2 540.9 16.8 21.4
-The results of correcting safety property violations on ACAS Xu are presented in Table 2.
-For the task of correcting safety property violation, AdvRepair, PRDNN, REASSURE, and ART successfully construct a provable repair, resulting in a 100% $\mathrm{RSR}$.
-In contrast, CARE attains an average $\mathrm{RSR}$ of 95.8%.
-Regarding $\mathrm{RGR}$, all baseline methods exhibit notable performance, with ART and our approach showcasing the highest performance at 100%.
-In terms of $\mathrm{FDD}$, our approach achieves the most favorable performance with zero, surpassing CARE, ART, PRDNN, and REASSURE.
-The results highlight the effectiveness of our approach in achieving provable repairs and maintaining high fidelity.
-Answer to RQ1: AdvRepair consistently outperforms the baselines in both local robustness repair and safety property violation correction tasks.
-It achieves a high repair success rate coupled with lower drawdown, showcasing superiority over the alternative approaches.
-The efficiency of AdvRepair is acceptable.
-4.3 Generalization
-In this experiment, we evaluate the generalization of the repaired network and its defense capabilities against new adversarial attacks, as shown in Table 3.
-AdvRepair outperforms the other baselines in terms of $\mathrm{RGR}$, showcasing superior performance.
-Notably, PRDNN, APRNN and TRADES also demonstrate some level of generalization, especially on the convolutional neural network, while the remaining tools exhibit relatively poor results.
-In terms of $\mathrm{DSR}$, AdvRepair exhibits superior performance compared to the other tools.
-TRADES is the only alternative showing a certain defensive capability against adversarial attacks on FNN_big and CNN, particularly when the radius is small.
-When evaluating $\mathrm{DGSR}$, we find that AdvRepair once again achieves the best performance.
-CARE, PRDNN, APRNN and TRADES demonstrate limited defense capabilities against FNN_big and CNN, especially when the radius is small.
-Notice that $\mathrm{DGSR}$ is the most significant metric for assessing generalization to global inputs.
-These results are sufficient to show that AdvRepair achieves a breakthrough in repairing adversarial attacks.
-Answer to RQ2: AdvRepair exhibits robustness defense against new adversarial attacks, surpassing the defensive performance of other tools.
-AdvRepair exhibits significant advantages in terms of the generalization of repairing adversarial attacks.
-Table 3: Results of generalization and defense against new adversarial attacks
+IV Experimental Evaluation
+In this section, we evaluate PatchPro by answering the following research questions:
+RQ1: What is the overall performance of PatchPro in repairing local robustness and correcting safety property violations?
+RQ2: Does the repaired DNN exhibit the capability to defend against new adversarial attacks?
+RQ3: Does PatchPro have scalability to repair large networks?
+RQ4: In the context of PatchPro, how does the size and quantity of patch modules influence the efficacy of the repair?
+IV-A Setup
+All the experiments are conducted on a machine with an AMD EPYC 7763 64-Core Processor, 256 GB of memory, and an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory.
+Each experiment has a timeout set to 10 000 seconds.
+Dataset
+We conduct evaluations on four common datasets: MNIST [29], CIFAR-10 [30], Tiny ImageNet[31], and ACAS Xu [32, 33].
+The first three are widely recognized benchmarks in the field for studying neural network robustness [34, 4], and ACAS Xu is a commonly used benchmark in research on neural network verification and repair [35, 22, 20].
+On MNIST, we use two fully connected networks and one convolutional network, while on CIFAR-10, we train a VGG19 [36] and a ResNet18 [37].
+On Tiny ImageNet, we train a Resnet152 and a WRN101-2 [38].
+We assess our approach on 35 ACAS Xu DNNs, which, as documented in [35], are expected to satisfy Property-2 but exhibit violations.
+For the local robustness repair task, we generate adversarial samples using PGD [4] for DNNs trained on MNIST and CIFAR-10 (50, 100, 200, 500, and 1,000 samples), as well as Tiny ImageNet (500 and 1,000 samples).
+The radius $r$ is set to 0.05, 0.1, and 0.3 for MNIST, $\frac{4}{255}$ and $\frac{8}{255}$ for CIFAR-10, and $\frac{2}{255}$ and $\frac{4}{255}$ for Tiny ImageNet.
+On ACAS Xu, we aim to repair the violation of Property-2 using one patch module while preserving the original performance.
+Although PatchPro repairs safety properties without requiring sample information, due to the needs of other tools such as CARE, we sample 500 counterexamples as the faulty inputs for repair.
+For CIFAR-10, the accuracies of VGG19 and ResNet18 are 93.4% and 88.3%, respectively.
+For Tiny ImageNet, the accuracies of WRN101-2 and ResNet152 are 64.4% and 68.2%, respectively.
+The maximum number of iterations $M$ in Alg.
+1 is set to 25.
+In Alg.
+2, the maximum number of epochs $R$, learning rate $\eta$, and selection number $K$ are set to 10, 10, and 800, respectively.
+All patch modules consist of a single linear layer, unless otherwise specified.
+On MNIST, the patch module takes the sample itself as input, while on CIFAR-10 and Tiny ImageNet, it takes the output from the network’s penultimate layer as input.
+For sampling the feature space, we use PGD attack with a step size of $\frac{2}{255}$ for 10 rounds, with each round consisting of 50 steps.
+We collect all adversarial examples generated during this process.
+Additionally, we use FGSM attack to generate 50 adversarial examples, which are then combined with the samples obtained from PGD.
+This combined set of samples is used for sampling the feature space.
+Baselines
+We compare PatchPro with the state-of-the-art repair methods including CARE [8], APRNN [22], PRDNN [20], REASSURE [21] and ART [25].
+Since APRNN and PRDNN require selecting a layer to repair on, we traverse all eligible layers across each baseline, and report the layer that exhibit the best performance.
+Additionally, for the local robustness repair task, we compare our approach with an adversarial training method TRADES [39], where we select the trained model with the best performance in 200 epochs with their default parameters.
+Metrics
+To assess generalization, we sample 10 adversarial examples on each $B(x_{i},r)$ different from $x_{i}^{*}$ to form a generalization set $D_{\mathrm{g}}$ for MNIST, CIFAR-10 and Tiny ImageNet, and use an independent set of 5 000 counterexamples for ACAS Xu.
+Besides these, we have an independent test set $D_{\mathrm{t}}$ of size 10 000, 10 000, 10 000 and 5 000 for MNIST, CIFAR-10, Tiny ImageNet and ACAS Xu, respectively.
+We employ AutoAttack [5] to attack the repaired DNN $F$ on $\{x_{i}\}_{i=1}^{n}$ and $D_{\mathrm{t}}$ with the same radius $r$.
+The metrics we use include repair success rate ($\mathrm{RSR}$), repair generalization rate ($\mathrm{RGR}$), drawdown ($\mathrm{DD}$), defense success rate ($\mathrm{DSR}$), and defense generalization success rate ($\mathrm{DGSR}$), defined as:
+$\mathrm{RSR}$ =|$=\frac{|\{\,i\mid C_{F}(x_{i}^{*})=C_{N}(x_{i})\,\}|}{n},$
+$\mathrm{RGR}$ =|$=\frac{|\{{\,x\in D_{\mathrm{g}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{g}}\right|},$
+$\mathrm{DD}$ =|$=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{N}(x)=\ell_{x}\,\}|-|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{t}}\right|},$
+$\mathrm{DSR}$ =|$=\frac{|\{\,i\mid\forall x\in\mathrm{AA}(F,B(x_{i},r)),C_{F}(x)=C_{N}(x_{i})\,\}|}{n},$
+$\mathrm{DGSR}$ =|$=\frac{|\{\,x\in D_{\mathrm{t}}\mid\forall x^{\prime}\in\mathrm{AA}(F,B(x,r)),C_{F}(x^{\prime})=C_{N}(x)\,\}|}{\left|D_{\mathrm{t}}\right|},$
+where $\ell_{x}$ denotes the ground truth of $x$, and $\mathrm{AA}(\varphi)$ represents the set of potential adversarial samples obtained by attacking the local robustness property $\varphi$ with AutoAttack.
+On ACAS Xu, the metric Drawdown is replaced with Fidelity Drawdown ($\mathrm{FDD}$) because there is no labelled ground truth for inputs on ACAS Xu:
+$\mathrm{FDD}=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)\neq C_{N}(x)\,\}|}{|D_{\mathrm{t}}|}.$
+The metrics $\mathrm{RSR}$ and $\mathrm{DD}$ (or $\mathrm{FDD}$) evaluates the overall performance of DNN repair by measuring the percentage of buggy inputs successfully repaired and how much the accuracy decreases, while $\mathrm{RGR}$, $\mathrm{DSR}$ and $\mathrm{DGSR}$ reflects how the repair generalizes to the robustness regions $B(x_{i},r)$ and to other inputs.
+TABLE III: Results of generalization and defense against new adversarial attacks
 Model $r$ $n$ RGR/DSR/DGSR/%
-CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
-FNN_ small 0.05 50 7.6 44.4 47.2 10.8 80.0 10.0 100.0 0.0 0.0 2.0 0.0 26.0 6.0 100.0 8.9 3.2 2.3 0.0 9.4 8.9 90.0
-100 2.2 48.7 46.7 9.1 0.0 10.0 100.0 1.0 0.0 0.0 0.0 2.0 20.0 100.0 9.3 5.1 0.8 0.0 9.8 11.3 96.4
+CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours
+FNN_ small 0.05 50 7.6 44.4 47.2 10.8 80.0 10.0 100.0 0.0 0.0 2.0 0.0 26.0 6.0 100.0 8.9 3.2 2.3 0.0 9.4 8.9 91.3
+100 2.2 48.7 46.7 9.1 0.0 10.0 100.0 1.0 0.0 0.0 0.0 2.0 20.0 100.0 9.3 5.1 0.8 0.0 9.8 11.3 96.6
 200 5.0 42.6 16.1 8.5 0.0 11.0 100.0 2.0 0.0 0.5 0.0 3.5 10.0 100.0 8.6 1.1 1.3 0.0 9.8 10.1 96.6
 500 6.6 39.0 10.3 7.2 0.0 11.0 100.0 1.4 0.0 3.8 0.0 2.8 9.6 100.0 8.3 0.5 5.3 0.0 9.8 8.9 96.6
-1000 12.7 41.3 10.9 2.0 11.0 100.0 2.3 0.0 0.4 3.9 14.4 100.0 6.1 0.2 0.9 0.0 9.8 11.3 96.6
+1000 12.7 41.3 10.9 2.0 11.0 100.0 2.3 0.0 0.4 3.9 14.4 100.0 6.1 0.2 0.9 9.8 11.3 96.6
 0.1 50 0.8 24.2 32.0 2.4 20.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 8.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.3
@@
 500 1.7 37.5 17.3 1.4 8.0 14.0 100.0 0.0 0.0 0.0 0.0 10.0 10.0 100.0 0.0 0.0 0.1 0.0 9.8 9.8 95.3
-1000 2.3 42.5 15.2 13.0 14.0 100.0 0.0 0.0 1.2 10.0 10.0 100.0 0.0 0.0 0.2 0.0 9.8 9.8 96.5
-0.3 50 0.0 24.2 37.4 0.8 20.0 8.0 94.8 0.0 0.0 0.0 0.0 8.0 8.0 92.0 0.0 0.0 0.0 0.0 9.8 9.7 86.4
-100 7.7 22.0 18.4 0.4 10.0 11.0 99.0 0.0 0.0 0.0 0.0 13.0 11.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.2
-200 0.0 26.1 8.8 0.5 5.0 8.0 99.5 0.0 0.0 0.0 0.0 10.5 9.5 100.0 0.0 0.0 0.0 0.0 9.8 9.6 90.8
-500 0.4 30.4 12.8 0.3 8.0 10.0 99.5 0.0 0.0 0.0 0.0 10.0 10.4 99.8 0.0 0.0 0.0 0.0 9.8 10.3 92.0
-1000 0.6 31.9 12.5 13.0 11.0 99.7 0.0 0.0 0.0 9.8 9.8 99.7 0.0 0.0 0.0 0.0 9.8 9.8 96.6
-FNN_ big 0.05 50 6.6 63.0 64.8 8.4 100.0 24.0 100.0 2.0 2.0 8.0 0.0 82.0 24.0 100.0 37.7 38.3 31.7 0.0 53.5 11.3 97.0
-100 10.9 63.3 67.4 9.6 90.0 9.0 100.0 7.0 1.0 11.0 0.0 60.0 9.0 100.0 35.7 39.8 19.0 0.0 54.9 10.3 97.1
-200 10.7 58.9 64.6 7.3 69.5 6.0 100.0 4.5 2.5 2.5 0.0 54.5 5.0 100.0 37.9 33.2 12.8 0.0 57.0 8.9 97.1
-500 10.6 54.8 66.2 78.8 10.0 100.0 4.8 0.6 0.0 52.8 7.4 100.0 35.4 29.7 1.7 0.0 61.2 10.1 97.2
-1000 11.8 54.6 68.4 79.2 20.0 100.0 7.1 0.4 0.0 56.1 19.2 100.0 36.8 25.5 1.8 0.0 66.2 11.3 97.2
-0.1 50 20.2 52.6 44.2 3.4 0.0 8.0 100.0 2.0 0.0 0.0 0.0 0.0 8.0 100.0 2.6 0.7 1.1 0.0 1.8 9.7 97.1
-100 2.0 47.5 46.3 3.2 20.0 11.0 100.0 0.0 0.0 0.0 0.0 7.0 11.0 100.0 1.8 0.9 0.0 0.0 4.0 9.8 97.1
-200 19.2 49.6 46.8 2.5 22.5 10.0 100.0 1.5 0.0 0.0 0.0 6.0 10.5 100.0 2.6 0.8 0.1 0.0 4.9 9.8 97.1
-500 1.2 50.6 49.3 23.6 15.0 100.0 0.0 0.2 0.0 2.8 9.0 100.0 1.2 0.2 0.0 0.0 5.4 8.9 97.2
-1000 1.0 51.1 49.9 42.1 11.0 100.0 0.0 0.0 0.0 7.2 10.5 100.0 1.4 0.1 0.0 0.0 9.2 10.3 97.2
-0.3 50 0.0 42.2 20.4 2.8 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.0
-100 0.1 43.4 27.8 1.4 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.7
-200 0.1 40.8 31.0 1.6 11.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.3 97.2
-500 0.3 36.9 36.0 9.2 9.0 100.0 0.0 0.0 0.0 0.0 10.2 100.0 0.0 0.0 0.0 0.0 0.0 10.3 97.2
-1000 0.4 36.5 37.1 15.0 12.0 100.0 0.0 0.0 0.0 0.0 11.8 99.9 0.0 0.0 0.0 0.0 0.0 11.3 97.2
-CNN 0.05 50 27.0 69.6 67.6 47.2 100.0 0.0 100.0 0.0 0.0 12.0 0.0 100.0 18.0 100.0 70.6 37.5 70.4 0.0 87.0 9.8 96.6
+1000 2.3 42.5 15.2 13.0 14.0 100.0 0.0 0.0 1.2 10.0 10.0 100.0 0.0 0.0 0.2 9.8 9.8 96.5
+0.3 50 0.0 24.2 37.4 0.8 20.0 8.0 100.0 0.0 0.0 0.0 0.0 8.0 8.0 100.0 0.0 0.0 0.0 0.0 9.8 9.7 96.5
+100 7.7 22.0 18.4 0.4 10.0 11.0 100.0 0.0 0.0 0.0 0.0 13.0 11.0 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.5
+200 0.0 26.1 8.8 0.5 5.0 8.0 100.0 0.0 0.0 0.0 0.0 10.5 9.5 100.0 0.0 0.0 0.0 0.0 9.8 9.6 96.6
+500 0.4 30.4 12.8 0.3 8.0 10.0 100.0 0.0 0.0 0.0 0.0 10.0 10.4 100.0 0.0 0.0 0.0 0.0 9.8 10.3 96.6
+1000 0.6 31.9 12.5 13.0 11.0 100.0 0.0 0.0 0.0 9.8 9.8 100.0 0.0 0.0 0.0 9.8 9.8 96.6
+FNN_ big 0.05 50 6.6 63.0 64.8 8.4 100.0 24.0 100.0 2.0 2.0 8.0 0.0 82.0 24.0 100.0 37.7 38.3 31.7 0.0 53.5 11.3 97.2
+100 10.9 63.3 67.4 9.6 90.0 9.0 100.0 7.0 1.0 11.0 0.0 60.0 9.0 100.0 35.7 39.8 19.0 0.0 54.9 10.3 97.2
+200 10.7 58.9 64.6 7.3 69.5 6.0 100.0 4.5 2.5 2.5 0.0 54.5 5.0 100.0 37.9 33.2 12.8 0.0 57.0 8.9 97.2
+500 10.6 54.8 66.2 78.8 10.0 100.0 4.8 0.6 0.0 52.8 7.4 100.0 35.4 29.7 1.7 61.2 10.1 97.2
+1000 11.8 54.6 68.4 79.2 20.0 100.0 7.1 0.4 0.0 56.1 19.2 100.0 36.8 25.5 1.8 66.2 11.3 97.2
+0.1 50 20.2 52.6 44.2 3.4 0.0 8.0 100.0 2.0 0.0 0.0 0.0 0.0 8.0 100.0 2.6 0.7 1.1 0.0 1.8 9.7 97.2
+100 2.0 47.5 46.3 3.2 20.0 11.0 100.0 0.0 0.0 0.0 0.0 7.0 11.0 100.0 1.8 0.9 0.0 0.0 4.0 9.8 97.2
+200 19.2 49.6 46.8 2.5 22.5 10.0 100.0 1.5 0.0 0.0 0.0 6.0 10.5 100.0 2.6 0.8 0.1 0.0 4.9 9.8 97.2
+500 1.2 50.6 49.3 23.6 15.0 100.0 0.0 0.2 0.0 2.8 9.0 100.0 1.2 0.2 0.0 5.4 8.9 97.2
+1000 1.0 51.1 49.9 42.1 11.0 100.0 0.0 0.0 0.0 7.2 10.5 100.0 1.4 0.1 0.0 9.2 10.3 97.2
+0.3 50 0.0 42.2 20.4 2.8 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.5
+100 0.1 43.4 27.8 1.4 2.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 14.0 100.0 0.0 0.0 0.0 0.0 0.0 11.3 96.5
+200 0.1 40.8 31.0 1.6 11.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 10.3 96.6
+500 0.3 36.9 36.0 9.2 9.0 100.0 0.0 0.0 0.0 0.0 10.2 100.0 0.0 0.0 0.0 0.0 10.3 97.2
+1000 0.4 36.5 37.1 15.0 12.0 100.0 0.0 0.0 0.0 0.0 11.8 100.0 0.0 0.0 0.0 0.0 11.3 97.2
+CNN 0.05 50 27.0 69.6 67.6 47.2 100.0 0.0 100.0 0.0 0.0 12.0 0.0 100. 18.0 100.0 70.6 37.5 70.4 0.0 87.0 9.8 96.6
 100 29.4 74.5 65.6 45.7 100.0 60.0 100.0 1.0 0.0 10.0 0.0 86.0 27.0 100.0 70.6 29.4 69.2 0.0 88.2 9.7 98.3
 200 27.6 79.8 62.5 50.9 97.0 0.0 100.0 1.0 0.0 3.0 0.0 84.0 20.5 100.0 70.6 10.4 70.6 0.0 89.9 9.7 98.3
-500 30.8 89.0 62.3 92.4 100.0 1.4 0.0 0.8 87.0 100.0 71.4 0.0 91.2 98.3
-1000 32.9 92.2 63.8 90.8 100.0 1.5 0.0 0.2 86.8 100.0 65.7 0.0 92.1 98.3
+500 30.8 89.0 62.3 92.4 100.0 1.4 0.0 0.8 87.0 100.0 70.6 4.3 71.4 91.2 98.3
+1000 32.9 92.2 63.8 90.8 100.0 1.5 0.0 0.2 86.8 100.0 70.7 0.9 65.7 92.1 98.3
 0.1 50 12.2 53.0 41.0 10.4 80.0 0.0 100.0 2.0 0.0 2.0 0.0 50.0 8.0 100.0 8.3 0.0 8.7 0.0 44.9 10.3 98.2
@@
 200 21.8 73.2 46.1 13.0 96.5 20.0 100.0 0.5 0.0 0.5 0.0 62.0 11.0 100.0 8.2 0.0 9.2 0.0 53.3 9.8 98.2
-500 30.4 84.5 47.0 93.2 100.0 0.2 0.0 0.0 66.6 99.8 8.4 0.0 62.4 98.2
-1000 24.6 90.3 48.7 94.7 100.0 0.6 0.0 0.0 70.1 99.9 4.6 0.0 67.8 98.3
+500 30.4 84.5 47.0 93.2 100.0 0.2 0.0 0.0 66.6 100.0 8.2 0.0 8.4 62.4 98.2
+1000 24.6 90.3 48.7 94.7 100.0 0.6 0.0 0.0 70.1 100.0 8.3 0.0 4.6 67.8 98.3
 0.3 50 0.0 44.8 70.0 4.0 0.0 10.0 100.0 0.0 0.0 0.0 0.0 0.0 12.0 100.0 0.0 0.0 0.0 0.0 0.0 9.8 98.3
@@
 200 0.0 69.4 78.2 3.6 12.5 0.0 100.0 0.0 0.0 0.0 0.0 0.0 7.0 100.0 0.0 0.0 0.0 0.0 0.0 8.9 98.3
-500 1.3 79.6 73.3 12.6 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.3
-1000 1.1 85.9 70.6 16.0 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.2
-4.4 Scalability
-We assess the scalability of AdvRepair on VGG19 and ResNet18 for local robustness repair using CIFAR-10 through the feature-space repair.
-The experimental results are presented in Table 4.
-Regarding $\mathrm{RSR}$, AdvRepair, PRDNN and APRNN achieve 100%.
-TRADES also performs commendably, although it slightly trails behind the provable tools.
-Notably, CARE exhibits an inability to repair the local robustness properties on the CIFAR-10 dataset.
-In terms of $\mathrm{RGR}$, AdvRepair demonstrates superior performance, with TRADES also achieving better results than other tools.
-As for $\mathrm{DD}$, AdvRepair outperforms the others except in scenarios where the number of buggy input is 1 000.
-In this situation, TRADES attains the best performance.
-This phenomenon is reasonable, since TRADES tends to perform better in the DNN accuracy when the number of samples in adversarial training are getting large.
-Looking at $\mathrm{DSR}$, we find that AdvRepair exhibits significantly better performance than the other tools, with the latter lacking the capability to defend against adversarial attacks on the original samples.
-For $\mathrm{DGSR}$, AdvRepair achieves the same level as the accuracy, surpassing the other tools.
-Although TRADES also performs comparatively well, it remains notably lower than AdvRepair.
-The running time of AdvRepair is acceptable, regarding the far superior repair performance and generalization.
-Answer to RQ3: AdvRepair demonstrates good scalability in repairing local robustness on large-scale DNNs, outperforming the other tools.
-Table 4: Results of repairing local robustness on CIFAR-10
+500 1.3 79.6 73.3 12.6 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0 98.3
+1000 1.1 85.9 70.6 16.0 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0 98.2
+IV-B Repair performance
+The results of the local robustness repair task on MNIST are summarized in Table I.
+We evaluate overall repair performance with $\mathrm{RSR}$, $\mathrm{DD}$, and runtime, representing fix success, accuracy retention, and speed, respectively.
+PRDNN, REASSURE, and PatchPro reach 100% $\mathrm{RSR}$ due to their provable designs.
+APRNN, despite being provable, underperforms on some FNN_small cases, while ART struggles with high-dimensional inputs.
+CARE and TRADES, lacking provable guarantees, show inferior $\mathrm{RSR}$ compared to provable methods.
+CARE’s suboptimal repair across all models may stem from the complex interplay between local robustness repair and DNNs’ numerous parameters, making it difficult to pinpoint and modify relevant parts.
+Although TRADES performs better on CNNs, its effectiveness wanes with larger-radius local robustness errors.
+In terms of $\mathrm{DD}$, PatchPro and REASSURE consistently outperform alternative methods across all models.
+The results of correcting safety property violations on ACAS Xu are presented in Table II.
+Each patch module here is of the same size as the original network.
+PatchPro, PRDNN, REASSURE, and ART achieve a perfect 100% $\mathrm{RSR}$ with provable repairs.
+Both ART and our method showcase remarkable $\mathrm{RGR}$ scores, achieving a perfect 100% among all the evaluated tools.
+PatchPro also excels with a $\mathrm{FDD}$ of zero, outperforming CARE, ART, PRDNN, and REASSURE.
+These results emphasize our method’s strength in delivering provable repairs and preserving high-fidelity functionality.
+TABLE IV: Results of repairing local robustness on CIFAR-10
 VGG19 ResNet-18
@@
 PRDNN 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
-TRADE 96.0 95.0 95.0 97.4 96.7 96.0 92.0 92.5 95.6 96.0 100.0 99.0 100.0 99.8 99.5 100.0 96.0 99.5 99.4 96.9
+TRADE 98.0 95.0 97.0 95.0 96.4 100.0 97.0 93.5 95.6 95.8 100.0 99.0 100.0 99.8 99.5 100.0 100.0 100.0 99.8 96.8
 APRNN 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
@@
 PRDNN 69.8 68.2 68.5 58.0 66.6 61.4 68.0 68.6 67.2 59.8 57.8 54.9
-TRADE 84.0 90.0 85.0 97.8 98.0 80.0 90.0 100.0 98.0 99.0 100.0 90.0 100.0 100.0 99.1 100.0 90.0 100.0 100.0 97.3
+TRADE 100.0 86.0 90.0 97.6 97.0 100.0 100.0 95.0 98.0 98.9 100.0 100.0 100.0 100.0 99.1 100.0 100.0 100.0 100.0 97.3
 APRNN 69.2 68.9 70.9 75.5 80.4 63.2 65.2 68.2 66.1 70.7 70.6 65.1 62.0 62.7 68.2 65.0 55.6 53.0 49.6 59.0
@@
 PRDNN 13.6 24.8 30.3 19.8 40.5 50.2 2.1 2.1 2.0 14.0 13.7 12.2
-TRADE 46.0 35.3 29.8 14.7 -3.3 43.9 37.8 29.7 11.8 -2.7 43.0 51.7 33.0 12.9 -11.2 50.4 51.4 29.2 13.7 -8.9
+TRADE 18.6 11.9 7.7 6.3 6.1 25.8 18.5 11.5 8.5 8.3 43.4 29.2 20.1 14.4 10.1 38.2 34.9 15.9 11.9 10.1
 APRNN 14.2 20.0 29.3 48.7 46.4 24.0 39.5 61.9 75.5 82.1 1.8 1.8 2.2 3.1 6.1 5.5 4.2 7.6 12.7 19.5
@@
 TRADE 3.6 3.7 2.9 3.4 3.4 4.1 3.7 3.2 2.3 2.6 5.7 6.7 6.0 5.4 5.2 5.6 5.0 5.4 5.8 5.2
-Ours 83.6 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3
+Ours 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3
 Time/s CARE 47.8 64.4 106.5 380.4 928.2 36.9 48.2 96.0 234.1 443.0 45.5 58.9 105.5 243.5 384.9 37.1 75.4 101.6 180.7 285.7
@@
 Ours 238.7 471.6 963.0 2516.9 5441.6 229.9 470.7 950.1 2485.0 5436.0 265.8 525.1 1050.3 2586.4 5262.3 264.7 523.1 1042.8 2627.8 5325.1
-4.5 Monotonicity w.r.t. the number of buggy behaviors
-We systematically evaluate the performance of the repair results concerning the number of buggy behaviors in the local robustness repair task.
-We consider five distinct configurations: 50, 100, 200, 500 and 1000.
-The comparative analysis encompasses the performance of various baselines alongside our approach across these configurations.
-AdvRepair not only showcases effectiveness in limiting the number of buggy behaviors but also excels when faced with a growing number of buggy behaviors.
-In contrast, APRNN encounters challenges in successfully repairing the FNN_small model when the number of buggy behaviors exceeds 100, and REASSURE experiences memory overflow issues when confronted with 1000 buggy behaviors.
-Answer to RQ4: Even in scenarios involving a large number of buggy behaviors, AdvRepair demonstrates efficient and effective removal of buggy behaviors while preserving the original performance of the models.
-Comparative analysis indicates that AdvRepair leverages the available buggy points to achieve superior repair results compared to alternative approaches.
-5 Related work
+IV-C Generalization
+In the experiment, we assess the repaired network’s generalization and resistance to new adversarial attacks (Table III).
+PatchPro leads in $\mathrm{RGR}$, demonstrating superior performance compared to other baselines.
+PRDNN, APRNN, and TRADES also show some degree of generalization, particularly on CNNs.
+For $\mathrm{DSR}$, PatchPro outperforms competitors, with TRADES showing moderate defense against small-radius attacks on FNN_big and CNN.
+Evaluating $\mathrm{DGSR}$, PatchPro again tops the list, while CARE, PRDNN, APRNN, and TRADES show limited defense against FNN_big and CNN.
+$\mathrm{DGSR}$ is key for gauging generalization to global inputs .
+Overall, these findings underscore PatchPro’ breakthrough in repairing adversarial attacks.
+Answer to RQ1 and RQ2: PatchPro consistently outperforms the baselines in both local robustness repair and safety property violation correction tasks.
+It achieves 100% repair success rate coupled with 0% drawdown and acceptable efficiency.
+Additionally, it demonstrates significant generalization against unforeseen adversarial attacks.
+IV-D Scalability Evaluation
+We examine PatchPro’ scalability on VGG19 and ResNet18 for CIFAR-10 local robustness repair via feature-space repair.
+Results in Table IV show PatchPro, PRDNN, and APRNN achieve 100% $\mathrm{RSR}$, while TRADES follows closely.
+PatchPro leads in $\mathrm{RGR}$ and $\mathrm{DD}$ with TRADES following suit.
+PatchPro also significantly outperforms others in $\mathrm{DSR}$ and matches accuracy levels in $\mathrm{DGSR}$, surpassing all tools.
+On larger datasets like Tiny ImageNet, PatchPro maintains superiority in repairing WRN101-2 and ResNet152, as seen in Table V.
+PatchPro’ running time is reasonable given its superior repair and generalization abilities.
+Although TRADES performs comparably, PatchPro consistently delivers higher performance across various metrics.
+For both datasets, patch modules are applied at the network’s second-to-last layer, using a fully connected network with a single linear layer.
+The linear layer takes the output from the network’s penultimate layer as input, and its output is added to the original network’s output.
+Answer to RQ3: PatchPro demonstrates good scalability in repairing local robustness on large-scale DNNs, which also achieves a 100% repair success rate coupled with 0% drawdown, outperforming the other tools.
+TABLE V: Results of repairing local robustness on Tiny ImageNet
+Model $r$ $n$ RSR/RGR/DD/DSR/DGSR/Time/s
+CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours CARE TRADE Ours
+WRN 101-2 $\frac{2}{255}$ 500 0.0 100.0 100.0 1.2 100.0 100.0 0.1 46.8 0.0 0.0 100.0 100.0 0.0 6.6 64.4 768.7 5040.8 2029.6
+1000 0.0 100.0 100.0 0.5 100.0 100.0 0.0 43.2 0.0 0.0 100.0 100.0 0.0 7.3 64.4 1345.7 7088.6 3260.7
+$\frac{4}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 51.8 0.0 0.0 97.6 100.0 0.0 1.8 64.4 721.9 4985.5 1731.8
+1000 0.0 100.0 100.0 0.0 100.0 100.0 0.1 49.5 0.0 0.0 97.2 100.0 0.0 2.1 64.4 1440.8 6976.2 3221.2
+ResNet 152 $\frac{2}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 64.5 0.0 0.0 100.0 100.0 0.0 5.2 68.2 545.3 4667.7 2219.6
+1000 0.0 100.0 100.0 0.0 100.0 100.0 0.0 62.6 0.0 0.0 99.9 100.0 0.0 6.1 68.2 1092.9 7503.5 4703.7
+$\frac{4}{255}$ 500 0.0 100.0 100.0 0.0 100.0 100.0 0.0 64.5 0.0 0.0 97.6 100.0 0.0 1.3 68.2 535.4 4624.2 2336.1
+1000 0.0 100.0 100.0 0.0 100.0 100.0 0.0 63.6 0.0 0.0 98.5 100.0 0.0 2.0 68.2 1028.1 7499.9 4366.8
+TABLE VI: The efficacy of patch module scale
+VGG19 ResNet-18
+$r=4/255$ $r=8/255$ $r=4/255$ $r=8/255$
+50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000
+DSR/PS 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
+PL 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.0 99.0 99.5 99.2 99.4 100.0 100.0 99.5 99.8 99.8
+DGSR/PS 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 93.4 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3 88.3
+PL 93.2 93.2 92.8 92.0 92.3 91.2 93.3 93.4 93.4 93.4 60.0 77.5 84.2 81.4 87.2 68.6 74.5 77.3 82.5 86.6
+Time/s PS 238.7 471.6 963.0 2516.9 5441.6 229.9 470.7 950.1 2485.0 5436.0 265.8 525.1 1050.3 2586.4 5262.3 264.7 523.1 1042.8 2627.8 5325.1
+PL 263.3 497.1 987.1 2562.7 4991.7 244.7 495.5 992.9 2468.7 5169.8 292.5 583.4 1155.1 3076.9 5865.6 341.5 592.4 1154 3007.7 5941.1
+Verified PS
+PL $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
+TABLE VII: The efficacy of single patch moudle
+VGG19 ResNet-18
+r = 4/255 r = 8/255 r = 4/255 r = 8/255
+50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000 50 100 200 500 1000
+RSR/10.0 18.0 16.5 18.2 16.7 10.0 9.0 14.5 11.8 11.6 12.0 18.0 16.5 16.2 16.2 10.0 8.0 12.0 11.6 11.3
+RGR/10.0 18.0 16.5 18.2 16.7 10.0 9.0 14.5 11.8 11.6 12.0 18.0 16.5 16.2 16.2 10.0 8.0 12.0 11.6 11.3
+DSR/6.0 4.0 2.5 1.8 1.3 6.0 6.0 3.0 2.0 1.3 4.0 1.0 3.5 1.8 1.3 8.0 6.0 4.0 4.8 4.1
+DGSR/5.8 5.2 3.9 3.9 3.8 5.7 5.6 2.05 1.9 1.9 5.2 1.5 1.5 1.5 1.4 7.9 7.2 5.4 4.8 4.8
+IV-E Impact of Patch Module Size and Quantity on Efficiency
+In this experiment, we evaluate the effectiveness of patch module size and quantity in repairing local robustness properties.
+We perform a comparative analysis of PatchPro’s performance across different sizes and quantities, presenting the experimental results in Table VI.
+The performance of PatchPro is assessed using VGG19 and ResNet18 on the CIFAR-10 dataset, across various combinations of patch module scales and quantities, including small patch modules (PS) and large patch modules (PL).
+Two distinct patch module sizes are considered: small and large.
+The small patch module consists of a single linear layer, which takes the output from the network’s penultimate layer as input.
+The large patch module also takes the output from the penultimate layer as input but additionally comprises three hidden layers with 200 neurons each.
+It is noted that larger patch networks tend to yield poorer results, possibly due to increased over approximation error when passing through ReLU nodes in the hidden layers.
+As the number of hidden layers increases, the cumulative over approximation error grows, leading to less effective repairs.
+Smaller patch scales slightly reduce the repair time.
+Additionally, when using DeepPoly to verify the feature layer repairs after applying smaller patches, the verification success rate is higher.
+We also tested the effect of using a single patch to repair the neural network, and the results are shown in Table VII.
+It can be observed that the repair effectiveness using a single patch is significantly inferior to using multiple patches.
+A single patch only achieves 10% to 20% RSR and RGR.
+In contrast, using multiple patches results in both RSR and RGR reaching 100%, as shown in Table IV.
+Additionally, the single patch performs much worse in terms of DSR and DGSR.
+Therefore, using multiple patches for repair is highly beneficial.
+Answer to RQ4: We observe that smaller patch modules are more effective, likely due to reduced over approximation error and shorter repair times, as well as higher verification success rates with abstract interpretation tools.
+For the aspect of patch module quantity, multiple patch modules significantly outperform only one patch module.
+V Related work
 Provable DNN repair.
-Our approach falls within the category of provable DNN repair methods.
 The methods most closely related to ours are ART [25] and REASSURE [21].
 In comparison to REASSURE, our approach fundamentally differs in its repair objectives.
-REASSURE focuses on fixing activation patterns of neural networks, while we directly guarantee the properties, whose input constraints may encompass a large number of activation patterns.
-In this way, our method exhibits superior generalization compared to REASSURE.
-Although ART also utilizes training in its repair process, it directly tunes parameters in the original neural network.
-In contrast, our approach involves training patch modules specifically for different properties, enhancing the specialization of our repairs and, consequently, their effectiveness.
+REASSURE focuses on fixing activation patterns of neural networks, yet for high-dimensional data, the input constraints of a property may encompass a large number of activation patterns.
+In contrast, PatchPro leverages formal verification to directly repair properties.
+Although ART also employs training in its repair process, it modifies parameters in the original neural network.
+On the other hand, PatchPro uses patch modules specifically to fix properties, ensuring the performance of the original network.
 Other methods such as PRDNN [20] and APRNN [22] formulate the repair problem in linear programming.
-However, their effectiveness diminishes in properties with high-dimensional polytopes, such as the robustness of image classifications.
+However, their efficacy is limited on properties with high-dimensional polytopes, such as the robustness of image classification.
 Heuristic DNN repair.
-Utilizing heuristic algorithms such as particle swarm optimization and differential evolution, CARE [61] and Arachne [59] aim to localize the neurons responsible for the fault.
-VeRe [41] focuses on providing formal verification guidance for fault localization and defining target intervals for repair synthesis.
-DeepRepair [82] and few-shot guided mix [51], for instance, expand the set of negative samples to generate additional training data.
-Tian et al.
-[64], on the other hand, augment the existing data by introducing real-world environmental effects, like fog, to the samples.
-To the best of our knowledge, DL2 [21] is the first paper to combine logical constraints and loss function, but it does not adopt convex approximation to give the loss function a certified guarantee.
+Utilizing heuristic algorithms such as particle swarm optimization and differential evolution, CARE [8] and Arachne [9] aim to pinpoint the neurons responsible for faults.
+VeRe [10] focuses on providing formal verification guidance to assist fault localization, and defines target intervals for repair synthesis.
+For example, DeepRepair [16] and few-shot guided mix [17] expand the set of negative samples to generate additional training data.
+Conversely, Tian et al.
+[40] augment the existing data by introducing real-world environmental effects, such as fog, to the samples.
+DL2 [24] fuses logical constraints and loss functions, but without convex-certified guarantees.
 DNN verification.
-In 2010, [50] proposed the first method for DNN verification.
-Over the past decade, various approaches of DNN verification has been proposed including techniques such as constraint solving [9, 25, 31, 38, 47, 33, 18, 28], abstract interpretation [24, 57, 80, 56, 36, 55], linear relaxation [30, 49, 79, 74, 7], global optimisation [52, 17, 53], CEGAR [2, 19, 48], reduction to two-player games [75, 77], and star-set abstraction [67, 66].
+Over the past decade, various approaches of DNN verification has been proposed including techniques such as constraint solving [41, 42, 43, 44], abstract interpretation [11, 45, 46, 47, 48], linear relaxation [49, 50, 51, 52, 53], global optimisation [54, 55, 56], CEGAR [57, 58, 59], reduction to two-player games [60, 61], and star-set abstraction [62, 63].
 These method offer provable estimations of DNN robustness.
@@
 Certified training [74, 75, 76] uses convex approximation in training the loss function, but it is only used to enhance the robustness of DNN, without making it accessible to repair properties in the training procedure.
-6 Conclusion
-We introduce AdvRepair, a novel approach for addressing adversarial attacks using limited data.
-Our method provides provable and targeted repairs within the robustness neighborhood, enabling the generalization of this defense to other inputs.
+VI Conclusion
+We introduce PatchPro, a novel approach for property-based repair of local robustness using limited data.
+Our method provides patch modules as neural networks to repair within the robustness neighborhood, enabling the generalization of this defense to other inputs.
 In terms of efficiency, scalability, and generalization, our approach surpasses existing methods.
-For future work, we plan to integrate BaB-based verification methods to enhance the performance of our approach, improving both precision and offering an alternative refinement approach.
-Additionally, our patch-based repair framework has the potential to evolve into a black-box repair method.
-Appendix 0.A Proof of Theorem 3.4
-Theorem 0.A.1
-Let $\varphi=(N,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
-$\mathcal{L}^{*}(\varphi):=\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}-r\cdot\bm{1})+\beta_{\ell},0),$
-where $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
-Proof
+VII Acknowledgements
+This work is supported by CAS Project for Young Scientists in Basic Research, Grant No.YSBR-040, ISCAS Basic Research ISCAS-JCZD-202302, and ISCAS New Cultivation Project ISCAS-PYFX-202201.
+The limitation of our work is that our patch allocation may assign inappropriate patches when the input is misclassified, and in this case PatchPro may fail to repair its local robustness.
+For future work, we plan to integrate verification methods based on Branch-and-Bound like [35] to enhance the performance of our approach, improving both precision and offering an alternative refinement approach.
+Additionally, our patch-based repair framework has the potential to evolve into a black-box repair method combined with black-box verification methods like [73].
+Appendix A Proof of Theorem 3.4
+Theorem 1.
+Let $\varphi=(F,B(x_{i},r))$ be a local robustness property. If $\mathcal{L}(\varphi)=0$ on $B(x_{i},r)$, i.e.,
+$\mathcal{L}^{*}(\varphi):=$ $\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x_{i}+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot$
+$(x_{i}-r\cdot\bm{1})+\beta_{\ell},0)=0,$
+where $\mathrm{elmax}$ and $\mathrm{elmin}$ are the element-wise $\max$ and $\min$ operation, $\bm{0}$ and $\bm{1}$ are the vector in $\mathbb{R}^{n_{0}}$ with all the entries 00 and $1$, respectively, then the property $\varphi$ holds.
+Proof.
 As $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ is a linear function with respect to $x$, we can calculate the maximum of $\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$ on ${B(x_{i},r)}$ as follows:
-$\max_{x\in B(x_{i},r)}\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$
-$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\beta_{\ell}$
-$=$ $\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\cdot(x_{i}+r)+\sum_{i\in\mathbb{R}^{n_{1}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\cdot(x_{i}-r)+\beta_{\ell}$
+$\quad\quad\max_{x\in B(x_{i},r)}\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}$
+$=\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x$
+$+\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\max_{x\in B(x_{i},r)}x+\beta_{\ell}$
+$=\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}>0\}}(\alpha_{\ell})_{i}\cdot(x_{i}+r)$
+$+\sum_{i\in\mathbb{R}^{n_{0}}}\mathbbm{1}_{\{(\alpha_{\ell})_{i}<0\}}(\alpha_{\ell})_{i}\cdot(x_{i}-r)+\beta_{\ell}$
 Then we have
 $\mathcal{L}(\varphi)(x)=\sum_{\ell\neq\ell_{0}}\max(\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell},0)=0,\forall x\in B(x_{i},r)$
-$\iff$ $\sum_{\ell\neq\ell_{0}}\max(\max(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x+r\cdot\bm{1})+\min(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x-r\cdot\bm{1})+\beta_{\ell},0)=0$
+$\iff$ $\sum_{\ell\neq\ell_{0}}\max(\mathrm{elmax}(\alpha_{\ell}^{\mathrm{T}},\bm{0})\cdot(x+r\cdot\bm{1})+\mathrm{elmin}(\alpha_{\ell}^{\mathrm{T}},\bm{0})$
+$\cdot(x-r\cdot\bm{1})+\beta_{\ell},0)=0$
 $\iff$ $\mathcal{L}^{*}(\varphi)=0.$
 Therefore, if $\mathcal{L}^{*}(\varphi)=0$, we have
-$\forall x\in B(x_{i},r),C_{N}(x)_{\ell}-C_{N}(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}\leq 0,$ (2)
+$\forall x\in B(x_{i},r),F(x)_{\ell}-F(x)_{\ell_{0}}\leq\alpha_{\ell}^{\mathrm{T}}x+\beta_{\ell}\leq 0,$
 then the property $\varphi$ holds. ∎
-Appendix 0.B Impact of Patch Module Size and Quantity on Efficiency
-In this experiment, we assess the efficacy of size and quantity of patch modules in the context of repairing local robustness properties.
-We conduct a comparative analysis of AdvRepair’s performance across various sizes and quantities, presenting the experimental outcomes in Table 5.
-Two distinct patch module sizes, namely small and large, are considered.
-The small patch module consists of three hidden layers with 64 neurons in the first two layers and 16 neurons in the last layer, while the large patch module comprises three hidden layers with 256 neurons in the first two layers and 64 neurons in the last layer.
-Additionally, we explore two different numbers of patch modules: property-guided patch modules and untargeted label patch modules.
-Table 5: The efficacy of patch module scale and quantity
-Model $r$ $n$ RSR/RGR/DSR/DGSR/%
-PS PL LS LL PS PL LS LL PS PL LS LL PS PL LS LL
-FNN_ small 0.05 50 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 90.0 90.0 89.9 89.8
-100 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.4 96.4 96.1 96.0
-200 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
-500 100.0 100.0 100.0 99.3 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.6
-0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.3 96.3 89.5 89.5
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.0 96.3 94.7 94.8
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.2 95.1 94.9 94.8
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 95.3 95.2 96.3 96.3
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.5 96.5 96.3 96.3
-0.3 50 94.0 94.0 94.0 94.0 94.8 94.8 94.8 94.8 92.0 92.0 92.0 92.0 86.4 79.9 44.6 44.5
-100 99.0 99.0 99.0 99.0 99.0 99.0 99.0 99.0 100.0 99.0 99.0 100.0 96.2 96.2 85.8 86.1
-200 99.5 99.5 99.5 99.5 99.5 99.5 99.5 99.5 100.0 99.0 97.5 98.0 90.8 91.2 85.9 85.9
-500 99.4 99.4 99.8 99.8 99.5 99.5 99.8 99.8 99.8 99.6 100.0 100.0 92.0 92.1 89.5 89.7
-1000 99.6 99.6 99.9 99.9 99.7 99.7 100.0 100.0 99.7 99.6 100.0 100.0 96.6 96.5 94.2 94.2
-FNN_ big 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.0 97.0 96.8 96.8
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.2
-0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.8 96.8
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.2 96.8 96.9
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.1 97.1 96.9 96.8
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 96.8 96.8
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 97.2 97.0
-0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.0 96.1 89.1 89.1
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.7 96.8 89.1 89.1
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 89.3 89.3
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 97.2 97.2 95.9 95.9
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.9 100.0 100.0 97.2 97.2 97.1 97.1
-CNN 0.05 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 96.6 96.6 96.6 96.7
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-0.1 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.0 98.0
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.1 98.0
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.5 99.5 99.5 98.2 98.2 98.0 98.1
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.8 99.8 99.8 99.8 98.2 98.2 98.0 98.1
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.9 99.8 99.8 99.8 98.3 98.3 98.0 98.0
-0.3 50 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-100 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-200 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-500 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.3 98.3 98.3 98.3
-1000 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.2 98.2 98.3 98.3
-The performance of AdvRepair is evaluated across different combinations of patch module scales and quantities, including small property-guided patch module (PS), large property-guided patch modules (PL), small untargeted label repairing patch modules (LS), and large untargeted label repairing patch modules (LL).
-Notably, the experimental findings indicate that both sizes of the patch modules exhibit commendable performance, with minimal impact observed on the repair performance due to the network size.
-Concerning the quantity of patch modules, the property-guided patch modules outperform the untargeted label repairing patch modules in defending against adversarial attacks within $D_{t}$.
-We hypothesize that this superiority is attributed to the specialization inherent in property-guided repair.
-Appendix 0.C Experimental Details
-Table 6: Network architectures and accuracies of DNN trained on MNIST dataset
+Appendix B An explainable example
+[…]
+Figure 2: A fully connected neural network $N$ with ReLU activations.
+Example 1.
+Consider the neural network in Fig.
+2 and the inputs $x=(-0.7,1)$ labeled “2”.
+Within the region $B(x,0.5)$, we have its counterexample $x^{*}=(-0.2,1.5)$, which violates the local robustness.
+To repair the network, we need to construct a patch $P=wx$ where $w\in\mathbb{R}^{2\times 2}$.
+After initializing $w$ to all $0.1$, we train the patch according the “while” loop begins at Line 7 in Alg.
+1.
+As shown in Line 5 in Alg.
+2), we first execute DeepPoly to obtain the safety violated loss function $\mathcal{L}=w_{1,1}x_{1}+w_{1,2}x_{2}-w_{2,1}x_{1}-w_{2,2}x_{2}+(0.7x_{1}+0.14x_{2}+1.08)$.
+To maximize $\mathcal{L}$, we set $x_{1}=-0.2$ and $x_{2}=1.5$, then we have $\mathcal{L}^{*}(w)=-0.2w_{1,1}+1.5w_{1,2}+0.2w_{2,1}-1.5w_{2,2}+1.15$.
+Finally, we update $w$ by $w-\eta\nabla\mathcal{L}^{*}(w)$ with the learning rate $\eta$.
+This process is repeated until the robustness is proven or the epochs reaches its limit $R$.
+To show the subsequent repair process, we set $\eta=0.6$ and $R=1$.
+Under this setting, the network is not be repaired, then we need to refine the robustness to two properties by bisecting the region $B(x,0.5)$ (see Line  14–19 in Alg.
+1).
+Specifically, by the judgment that $\partial_{1}\mathcal{L}>\partial_{2}\mathcal{L}$, we select the dimension of $x_{1}$ and divide the robustness region $B(x,0.5)$ into $[-1.2,-0.7]\times[0.5,1.5]$ and $[-0.7,-0.2]\times[0.5,1.5]$.
+We perform the above repair process again, and based on a more accurate abstraction provided by two new properties, we finally obtain the repaired network with the patch
+$P=\begin{pmatrix}0.22&-0.8\\ -0.02&1\end{pmatrix}\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}\,.$
+Appendix C Limitation
+As mentioned in section III-C, the adversarial examples not within any known robustness regions may lead to the establishment of incorrect new properties, and the resulting inappropriate patches allocation may ultimately affect the performance of the repaired neural network.
+Although this situation is rare in reality, we created an extreme setting here to investigate this weakness of our method.
+Specifically, considering the local robustness with radius $r=4/255$, we reuse the repaired VGG19 for CIFAR-10 and execute it over the adversarial dataset $D_{\mathrm{adv}}$ consisting of the adversarial examples generated by attacking the original network over the testset $D_{\mathrm{t}}$.
+The attack utilize AutoAttack with the step size of 10.
+By preserving the new properties established in this execution, we retest the drawdown of the repaired model over $D_{\mathrm{t}}$.
+The results are presented in Fig.
+3, showing that the accuracy of the repaired model drops significantly under this extreme setting.
+This is a limitation of PatchPro, and it also echoes the statement aforementioned that adversarial detection can serve as an important supplement to our method.
+Figure 3: Results under the extreme setting.
+Appendix D Network architectures and accuracies of the DNN trained on MNIST
+The network architectures and accuracies of the DNN trained on MNIST are detailed in Table VIII.
+TABLE VIII: Network architectures and accuracies of DNN trained on MNIST dataset
 Name Accuracy/Model structure
@@
 linear layer of 10 hidden units
-FNN_big 98.3 linear layer of 200 hidden units
+FNN_big 97.2 linear layer of 200 hidden units
 linear layer of 200 hidden units
@@
 linear layer of 10 hidden units
-CNN 97.2 Conv2d(1, 16, 4, stride=2, padding=1)
+CNN 98.3 Conv2d(1, 16, 4, stride=2, padding=1)
 linear layer of 100 hidden units
 linear layer of 10 hidden units
-The network architectures and accuracies of the DNN trained on MNIST are detailed in Table 6.
-For CIFAR-10, the accuracies of VGG19 and ResNet18 are 93.4% and 88.3%, respectively.
-In Algorithm 1, the maximum number of iterations $M$ is set to 25.
-In Algorithm 2, the maximum number of epochs $R$, learning rate $\eta$, and selection number $K$ are set to 10, 10, and 800, respectively.
-The attack methods in AutoAttack have a maximum of 10 iterations, and the number of restarts is set to 1.